### v1

# Sparse Portfolio Selection via Topological Data Analysis based Clustering

Abstract
This paper uses topological data analysis (TDA) tools and introduces a data-driven clustering-based stock selection strategy tailored for sparse portfolio construction. Our asset selection strategy exploits the topological features of stock price movements to select a subset of topologically similar (different) assets for a sparse index tracking (Markowitz) portfolio. We introduce new distance measures, which serve as an input to the clustering algorithm, on the space of persistence diagrams and landscapes that consider the time component of a time series. We conduct an empirical analysis on the S&P index from 2009 to 2020, including a study on the COVID-19 data to validate the robustness of our methodology. Our strategy to integrate TDA with the clustering algorithm significantly enhanced the performance of sparse portfolios across various performance measures in diverse market scenarios.
keywords:
Topological Data Analysis, Clustering, Index Tracking, Mean-Variance Portfolio, Global Minimum Variance Portfolio, Sparse Portfolios
journal: Decision Support Systems
1 Introduction
Asset selection and portfolio optimization are critical tasks for financial institutions such as banks, fund management firms, and individual investors as they seek to manage risk and maximize returns in an increasingly competitive business environment and complex market conditions. Over the years, academics and practitioners have actively developed quantitative techniques and models to assess risk and construct efficient portfolios that perform well out-of-sample. In this pursuit, sparse portfolio selection has become a well-known and extensively researched problem in the fields of financial economics and optimization, where sparseness is a desirable property not only from the point of view of transaction costs but also because it allows investors to concentrate their capital on assets that are most likely to generate returns while minimizing exposure to unforeseen risks or market fluctuations. Moreover, it also offers a potential solution to the limitations of traditional mean-variance or index-tracking portfolios, as it focuses on a smaller subset of assets to improve portfolio performance. [^1] Despite its popularity, mean-variance analysis has received criticism for its susceptibility to parameter estimates and assumptions. It is observed that small changes in assumed asset returns, volatilities, or correlations lead to substantial effects on the optimization results, ultimately leading to sub-optimal performance when tested against real-world data. In essence, the classic Markowitz portfolio optimization represents an ill-posed inverse problem. Several techniques have been proposed to address this challenge, including regularization techniques imposing sparseness and adopting shrinkage estimators of expected returns and covariance matrix. [^2] In index tracking, the goal is to construct a portfolio of securities that replicates the performance of a market index. The simplest method to achieve this goal is through full replication, which entails holding all the assets in the target market index in the same proportions. While full replication provides perfect tracking performance in a frictionless market, it is expensive, requiring frequent re-balancing of many stocks. It is sometimes not possible due to restrictions on trading illiquid securities or managerial constraints. An alternative approach is partial replication, which aims to mimic the index performance using a small subset of securities.
The problem of sparse portfolio selection can be reformulated as a two-step problem, with the first step of selecting a subset of securities based on their risk and return characteristics and the second step of obtaining the optimal weights. In this paper, we adopt clustering to select a smaller set of securities that provide diversification benefits to the portfolio. Clustering methods have emerged as one of the most innovative strategies in portfolio selection in recent years. This technique for classifying data objects into homogeneous groups based on their similarities proved to be effective in constructing sparse portfolios (Panton et al. (1976); Hui (2005)).
This two-step approach has several advantages over the alternative methods in the literature, such as regularization techniques (imposing a constraint on the $\ell_{q}$-norm of portfolio weights) or including a cardinality constraint on the number of securities to be included in the portfolio to promote sparsity and stable allocations. Adding a cardinality constraint significantly increases the combinatorial complexity of the optimization problem, resulting in a mixed-integer quadratic programming problem that is generally NP-hard and computationally intractable. Heuristic methods, such as genetic algorithms, are typically required to find approximate solutions, which may not guarantee global optimality (Ruiz-Torrubiano and Suárez (2009); Woodside-Oriakhi et al. (2011); Fastrich et al. (2014)). Similarly, the regularization technique involves determining the regularization coefficients through cross-validation, which becomes particularly challenging when each security has its regularization parameter. This approach has been shown to outperform the approach with a common regularization parameter. [^3] For example, Brodie et al. (2009) suggests integrating an $\ell_{1}$-norm (LASSO) penalty on the vector of portfolio weights in optimization models to regularize the optimization problem and encourage sparse portfolios. Moreover, DeMiguel et al. (2009a) demonstrate that norm-constrained global minimum variance portfolio yields improved out-of-sample performance and Giamouridis and Paterlini (2010) and Giuzio et al. (2018) utilize $\ell_{1}$ penalty in the context of replicating hedge funds. Recently, Ho et al. (2015) considered the mean-variance optimization with an advanced regularization technique called weighted Elastic-Net. Further, these methods tend to result in risk concentration as the resulting portfolio may not be diversified. This can lead to over-exposure to a single industry sector, thus making the portfolio riskier due to vulnerability to a downturn in that sector. Furthermore, due to local optima and discontinuities in the search space, both the cardinality constraint problem and the $\ell_{q}$-norm constraint problem are difficult to solve from an optimization standpoint. Finally, the regularization technique LASSO suffers from several known shortcomings, such as selecting at random from equally correlated assets (Bondell and Reich (2008)), being biased for large coefficient values (Fan and Li (2001); Giuzio et al. (2018)), and being ineffective in the no-short-sale area (DeMiguel et al. (2009a)). In contrast, clustering can result in a more diversified and sparse portfolio without increasing the computational burden. Moreover, clustering allows for the combination of judgment and quantitative optimization as it is an intuitive approach.
The key to any clustering technique is the distance function that enables it to identify similarities and dissimilarities among the inputs in the data set (time series of stock returns in our case) and allocate them to different clusters. Traditionally, correlation-based clustering has been employed to locate clusters in the portfolio of financial assets, for instance, Mantegna (1999). However, it is well known that Pearson’s correlation is not a suitable measure of dependence to be used outside the Gaussian (elliptical) distributions, and, in particular, it does not give an accurate indication and understanding of the actual dependence between risk exposures Embrechts et al. (2003). Furthermore, even if two assets are perfectly correlated, an investment in one can lead to positive and negative returns in the other. In this paper, we employ a recently introduced approach known as Topological Data Analysis (TDA) to define the similarity between two time series and propose an efficient data-driven approach to sparse portfolio selection. The underlying idea of TDA is that persistent homology, through filtration, enables highly flexible and multi-resolution analysis of $n$-dimensional holes, which are the most salient features of a data set. TDA thus allows us to relate objects based on their qualitative structural properties, such as shape and connectedness. This paper argues that topological features can help identify exciting time series patterns and spatial data clustering. In particular, insights can be gained by examining the essential qualitative characteristics of data sets that are often overlooked by traditional distance-based clustering techniques. To our knowledge, this paper is the first to present a pioneering application of TDA to address the sparse portfolio selection using clustering.
Our paper makes several methodological and empirical contributions that address the aforementioned issues. First, despite the considerable amount of research that has been conducted on applying TDA to study financial time-series, there still needs to be practical implementation and utilization of this method in the field of quantitative finance. [^4] Recently, Goel et al (2023) Goel et al. (2023) proposed TDA based sparse index tracking model where the authors adopt the regularization approach to achieve sparseness. We leverage tools from TDA, namely persistence diagrams and landscapes, and design strategies for sparse portfolio selection. By doing so, we establish a link between the growing research on TDA and the most frequently studied portfolio selection problem in the financial economics literature. Thus, we contribute to the emerging literature on applications of TDA to the problem of selecting sparse portfolios.
Second, our research draws attention to a critical limitation of distance metrics defined on persistence diagrams when applied to financial time-series data. Specifically, we show that the Wasserstein distance between the PDs corresponding to the return series of two hypothetical assets is zero, even when the two return series are not identical. We identify the reason for this behavior as the critical time component in financial data. To address this issue, we introduce two novel distance measures. The first, called the differenced Wasserstein distance (DWD), defines the distance between the persistence diagrams of two time series, ${x_{t}}$ and ${y_{t}}$, as the Wasserstein distance between the persistence diagram of the differenced time series ${x_{t}-y_{t}}$ and that of a constant series. The second called the average Wasserstein distance (AWD), partitions the two time series into overlapping sub-series and calculates the average Wasserstein distance between the corresponding sub-series. We also introduce similar measures, including the average landscape distance (ALD) and the differenced landscape distance (DLD) on the space of persistence landscapes. Moreover, we also demonstrate the stability of the persistence diagrams concerning the newly introduced distance measures, thus contributing to the literature on TDA.
Third, we utilize the introduced distance measures to propose two strategies for sparse portfolio selection. Our first strategy focuses on constructing a sparse index tracking (IT) portfolio, while our second strategy aims to select sparse mean-variance (MV) and global minimum variance (GMV) portfolios. To accomplish this, we create a distance matrix corresponding to the index and its $n$ constituents and define the corresponding similarity matrix. We then use this matrix as input for affinity propagation clustering (APC). We select the assets that share the cluster with the index for the sparse IT portfolio, which are topologically similar to the index by construction. We solve an index tracking problem on the selected assets. We select representative assets from each cluster for the sparse MV and GMV portfolios and solve the portfolio optimization problem on the filtered assets, thus obtaining a sparse portfolio. Our TDA-based similarity matrices and the APC offer an entirely data-driven approach, unlike the methodologies that employ $k$-mean clustering for sparse portfolio selection, which requires investors to pre-specify the number of assets in the portfolio.
We conduct an extensive empirical analysis, investigating the performance of the sparse index tracking, MV, and GMV portfolios obtained using the daily returns of the S&P 500 index and its constituents spanning approximately a decade from Dec 2009 to Dec 2019. We further assess the performance of our strategies during the recent COVID-19 Pandemic period from July 2019 to November 2020 to check the robustness. We also compare the performance of our sparse index tracking model on the benchmark data set for index tracking problems available publicly on the Beasley OR library. Our analysis shows that our TDA-based sparse portfolios outperform the traditional correlation-based sparse portfolios as well as other leading benchmark models in the literature.
The rest of this paper is structured as follows. Section 2 briefly reviews TDA. Section 3 introduces the novel distance measures and demonstrates their stability. Section 4 describes the methodology, starting with the problem of sparse portfolio selection, followed by a brief review of affinity propagation clustering, and finally, our portfolio selection strategies. Section 5 presents the data and the empirical findings. Section 6 concludes with potential future research directions.
2 Background of TDA
Inspecting the 2-dimensional point clouds in Figure 1, it is evident that each point cloud has a distinct shape. [^5] A point cloud is a finite set of points in some Euclidean space $\mathbb{R}^{d}$ (or a more general metric space). Specifically, the points in the first point cloud are scattered around one loop, while those in the second point cloud are dispersed around two loops. Given our inability to visualize beyond two or three dimensions, the natural question is how to formalize these qualitative observations, particularly the shapes of point clouds in higher dimensions. Fortunately, TDA provides a rigorous tool to systematically define and study the “shape” of data in high dimensional spaces.
Figure 1: Examples of two-dimensional noisy point clouds.
The first step involves building a simplicial complex given the point cloud $X=\{x_{i}\}^{N}_{i=0}$ in $\mathbb{R}^{q}$. A simplicial complex is a space comprising points, edges, triangles, tetrahedra, and higher-dimensional polytopes. In the abstract sense, a simplicial complex is a finite set $S$ of non-empty subsets of a set $S_{0}$ such that $\{v\}\in S$ for every $v\in S_{0}$, the intersection of any two simplices in $S$ is either empty or shares faces, and $\tau\subset\sigma$ and $\sigma\in S$ ensure that $\tau\in S$ [^6] A $k$-simplex is a convex hull $[a_{0},a_{1},\ldots,a_{k}]=\{\sum_{i=0}^{k}\eta_{i}a_{i}:\sum_{i=0}^{k}\eta_{i}=1\}$ of $k+1$ geometrically independent points $\{a_{0},a_{1},\ldots,a_{k}\}$. Furthermore, the faces of a $k$-simplex $[a_{0},\ldots,a_{k}]$ are the $(k-1)$-simplices spanned by subsets of $\{a_{0},\ldots,a_{k}\}$. Note that 0-dimensional simplices represent data points (vertices), 1-dimensional simplices are connected pairs of vertices (edges), and a 2-dimensional simplex is a filled triangle determined by its three vertices. There are several ways to construct a simplicial complex for a given point cloud, including the Cech complex and the Vietoris-Rips complex. We adopt the Vietoris-Rips complex in this paper. The underlying idea of the Rips complex is to connect the pairs that are sufficiently close to points (vertices) by edges. Mathematically,
Definition 2.1.
Given a resolution parameter $\epsilon>0$, the Vietoris–Rips complex of $X$ is defined to be the simplicial complex ${\mathcal{R}}_{\epsilon}(X)$ satisfying $[x_{i1},\ldots,x_{il}]\in{\mathcal{R}}_{\epsilon}(X)$ if and only if $diam(x_{i1},\ldots,x_{il})<\epsilon$.
The next question is how to extract the “shape” of the point cloud from the simplicial complex. TDA defines this shape by examining the topological features present in this complex. More specifically, TDA extracts the topological features characterized by $k$-dimensional holes, where $k=0,1,2$ represents connected components, holes, and voids, respectively, and uses these features as a surrogate to describe the shape of the point cloud.
Another important question revolves around the choice of $\epsilon$, i.e., how do we define “sufficiently close” points to construct a simplicial complex? The choice of $\epsilon$ is vital as a very large value can lead to just one connected component as all the points get connected to each other, hence a loss of any meaningful topological features. On the other hand, we might not see any interesting topological features when $\epsilon$ is too small due to no edge formation. TDA offers a systematic solution to address this question. Instead of making random choices of $\epsilon$ that could impact our conclusions, TDA involves analyzing the shape of the point cloud across various resolutions by studying the evolution of topological features as a function of $\epsilon$. To accomplish this, we construct a sequence of Vietoris–Rips simplicial complexes for $0=\epsilon_{0}<\epsilon_{1}<\ldots<\epsilon_{n}$ with $\{\epsilon_{n}\}\in\mathbb{R}^{+}\cup\{0\}$. We call this sequence of complexes, a Vietoris–Rips filtration and denote it by $\{{\mathcal{R}}_{\epsilon_{n}}(X)\}_{n\in\mathbb{N}}$. Within this filtration, due to several topological transformations, such as the closure of existing holes or the emergence of new edges, several topological features would appear (birth) and disappear (death) in the corresponding simplicial complex. We finally assign each topological feature a ‘birth’ and ‘death’ value. The persistence (age) of these features provides a meaningful characterization of the overall shape of the data; the features with longer persistence are considered more robust, significant, and representative of the shape of the point cloud. Conversely, features with shorter persistence are considered less significant or noise and can be filtered out if necessary.
Figures 2 depict this step-by-step procedure, also known as persistence homology, for obtaining the filtration from the point cloud. The idea is to encircle each point in the point cloud with $\epsilon$-sized balls, gradually expanding $\epsilon$ until all the balls merge, resulting in one connected component and keeping track of birth and death times of various topological features, such as connected components, loops, and voids.
(a) $\epsilon_{1}$=0.2, loops=0
(b) $\epsilon_{3}$=0.6, loops=1
(c) $\epsilon_{5}$=1.5, loops=2
(d) $\epsilon_{6}$=2.5, loops=0
Figure 2: A diagram depicting the Rips filtration process. Each point in the point cloud is equipped with an equal-sized ball, with the radius of the ball serving as the filtering parameter. By increasing the value of the radius, a succession of nested simplicial complexes is formed. It causes characteristics like connected components and holes to emerge and vanish. The initial loop appeared in 2(b) and died in 2(c). In 2(c), two additional substantial large holes can be observed, which die in 2(d).
Having obtained the Rips filtration, we summarize the birth and death times of topological features in a persistence diagram (PD) ${\mathcal{D}}$, which is a multi-set of points in $\mathbb{W}\times\{0,1,\ldots,q-1\}$, where $\mathbb{W}:=\{(b,d)\in\mathbb{R}^{2}:d\geq b\geq 0\}$ and each element $(b,d,f)$ represents a topological feature of dimension $f$ that appeared at $\epsilon=b$ and disappeared at scale $\epsilon=d$ during the construction of filtration. Figure 3(a) gives the persistence diagram corresponding to the filtration in Figure 2. When we study persistence diagrams for data analysis, one must be able to compare persistence diagrams. Therefore, it becomes necessary to introduce a distance metric between two persistence diagrams to describe their relation. The similarity between two persistence diagrams is commonly measured by $p$-Wasserstein distance, defined as
Definition 2.2.
The $p$-Wasserstein distance Cohen-Steiner et al. (2010) between two persistence diagrams ${\mathcal{D}}_{{X}}$ and ${\mathcal{D}}_{{Y}}$ corresponding to data sets $X$ and $Y$ is defined as
$WD_{p}({\mathcal{D}}_{{X}},{\mathcal{D}}_{{Y}})=\left(\inf_{\gamma}\sum_{a\in{{\mathcal{D}}_{{X}}}}||a-\gamma(a)||_{\infty}^{p}\right)^{\frac{1}{p}}.$
where $p\geq 1$ and $\gamma$ ranges over all bijections from the points in ${\mathcal{D}}_{{X}}$ to the points in ${\mathcal{D}}_{{Y}}$. For $p=\infty$, $p$-Wasserstein distance is called the bottleneck distance.
Though persistence diagrams contain potentially valuable information about the underlying data set, it is a multi-set which, when equipped with Wasserstein distance, forms an incomplete metric space and hence is not appropriate to apply tools from statistics and machine learning for processing topological features (Bubenik et al. (2015)). A persistence landscape is an alternative tool to convey the information in the persistence diagram. It comprises a sequence of continuous and piece-wise linear functions defined on a re-scaled birth-death coordinate, with the peaks of a persistence landscape representing the significant topological features. Given the birth-death pairs $p_{i}(a_{i},b_{i}),~{}i=1,\ldots,m$, a persistence landscape, illustrated in Figure 3(b), is defined as a sequence of functions $\lambda:\mathbb{N}\times\mathbb{R}\rightarrow[0,\infty)$, denoted as $\lambda(k,t)=\lambda_{k}(t)$. Here, $\lambda_{k}(t)$ represents the $k$-th largest value among the set $\{\Lambda_{p_{i}}(t)\mid i=1,\ldots,m\}$ and we set $\lambda_{k}(x)=0$, ensuring that $\lambda_{k}(t)=0$ for $k>m$ if the $k$-th largest value does not exist. Here, for a fixed feature dimension $f$, we associate a piece-wise linear function $\Lambda_{p}(t):\mathbb{R}\rightarrow[0,\infty)$ with each birth-death pair $p=(a,b)\in{\mathcal{D}}$, where ${\mathcal{D}}$ is the persistence diagram, as follows:
$\Lambda_{p}(t)=\left\{\begin{array}[]{ll}t-a&\quad t\in[a,\frac{a+b}{2}]\\[5.0pt] b-t&\quad t\in[\frac{a+b}{2},b]\\[5.0pt] 0&\quad\mbox{otherwise}.\end{array}\right.$ (4)
Simply put, a persistence landscape is obtained by rotating the persistence diagram 45 degrees clockwise and then drawing right-angle isosceles triangles, treating the birth-death pair in the persistence diagram as a right-angle vertex. Finally, individual functions are traced out from this collection of right triangles, such that the first ($k$th) landscape function is the point-wise ($k$th) maximum of all the triangles drawn.
One of the notable advantages of the persistence landscape representation, as compared to persistence diagrams (refer to Bubenik et al. (2015); Bubenik (2018)), is that it forms a subset of the Banach space $L^{p}(\mathbb{N}\times\mathbb{R})$, which comprises sequences $\lambda=(\lambda_{k})_{k\in\mathbb{N}}$. This set possesses an inherent vector space structure and becomes a Banach space when equipped with a norm that measures the degree of “spread out” or “concentration” exhibited by the landscape functions. The norm of a persistence landscape characterizes a point cloud by a single numerical value, and a higher norm value indicates the presence of more significant topological features in the underlying point cloud and is given by
$||\lambda||_{p}=\left(\displaystyle\sum_{k=1}^{\infty}||\lambda_{k}||^{p}_{p}\right)^{\frac{1}{p}},$ (5)
where $||\cdot||_{p}$ is the $L^{p}$-norm.
(a) Persistence diagram
(b) Persistence Landscape
Figure 3:   (a) The persistence diagram associated to Figure 2; the dots in it represent the birth and death of a feature; for instance, three off-diagonal red dots represent the birth and death of three significant holes in Figure 2;  (b) the corresponding persistence landscape.
Given that TDA operates on point clouds in multi-dimensional spaces, and our focus in this paper is on time series of stock returns, which lack natural point cloud representations, we must devise a method to obtain a point cloud from a time series. To achieve this, we adopt the standard procedure, called Taken’s embedding (Takens (1981)) that embeds a time series $x=\{x_{t},t=1,\ldots,T\}$ to a multi-dimensional representation as follows:
(a) Time series
(b) Point cloud
Figure 4: Time series and point cloud in $\mathbb{R}^{2}$ constructed using Takens’ embedding
${X}=\left[\begin{array}[]{c}X_{1}\\ X_{2}\\ \vdots\\ X_{T-(d-1)\tau}\\ \end{array}\right]=\left[\begin{array}[]{cccc}x_{1}&x_{1+\tau}\dots&x_{1+(d-1)\tau}\\ x_{2}&x_{2+\tau}\dots&x_{2+(d-1)\tau}\\ \vdots&\vdots&\vdots\\ x_{T-(d-1)\tau}&x_{T-(d-2)\tau}\dots&x_{T}\\ \end{array}\right],$ (6)
where $d$ is the dimension of reconstructed space (embedding dimension), $\tau$ is the time delay, and $T-(d-1)\tau$ is the number of points in the new space. The row of the matrix $X$ denotes a point in the new space. We present the idea of Taken’s embedding in Figure 4, where the first plot shows the time series. In contrast, the second plot presents the corresponding 2-dimensional point cloud, obtained using Taken’s embedding with $d=2$ and $\tau=1$.
3 New distance measures
The notion of distance is inherently linked to the characteristics of the underlying data. For instance, the distance between two financial time series may differ from that between two DNA sequences. Nonetheless, distance measure aims to capture some sense of similarity, and often, one can employ the distance measure inherited from the reference space. However, if that is not the case, one must use a more suitable metric. For example, to compare persistence diagrams, Wasserstein distance (or bottleneck distance) is traditionally used due to their stability concerning perturbations of the input Gidea and Katz (2018); Gidea (2017); Goel et al. (2020). However, considering the time factor while dealing with financial time series is crucial, which the traditional Wasserstein distance fails to capture, as shown in the following toy example.
Toy example 1 Suppose we have two time series obtained from some experiment, ${x}=\{x_{t},~{}t=1,\ldots,T\}$ and ${y}=\{y_{t},~{}t=1,\ldots,T\}$ in $\mathbb{R}^{T}$ such that ${y}$ is the reverse of ${x}$, that is, if ${x}=(x_{1},x_{2},\ldots,x_{T})$ then ${y}=(x_{T},x_{T-1},\ldots,x_{1})$. The point clouds obtained using Takens’ embedding theorem for $x$ and $y$ (say for $d=3$ and $\tau=2$) are
${X=\left[\begin{array}[]{ccc}x_{1}&x_{3}&x_{5}\\ x_{2}&x_{4}&x_{6}\\ \vdots&\vdots&\vdots\\ x_{T-4}&x_{T-2}&x_{T}\\ \end{array}\right],Y=\left[\begin{array}[]{ccc}y_{1}&y_{3}&y_{5}\\ y_{2}&y_{4}&y_{6}\\ \vdots&\vdots&\vdots\\ y_{T-4}&y_{T-2}&y_{T}\\ \end{array}\right]=\left[\begin{array}[]{ccc}x_{T}&x_{T-2}&x_{T-4}\\ x_{2}&x_{4}&x_{6}\\ \vdots&\vdots&\vdots\\ x_{5}&x_{3}&x_{1}\\ \end{array}\right].}$
since $y$ is reverse of $x$. Since TDA is a coordinate-free approach (or coordinate invariant), the point clouds for $x$ and $y$ will result in the same persistence diagrams and hence zero Wasserstein distance, implying a perfect similarity between the two, which is misleading.
While the example above is hypothetical, it advocates for a more appropriate distance measure for time-series data. It emphasizes that Wasserstein distance might be beneficial when dealing with observations from a fixed geometrical object where one is trying to identify the structure of that object. Still, it is unsuitable for financial applications where time is essential. Further, the literature has observed that computing the Bottleneck or the Wasserstein distances is computationally demanding as it requires finding point correspondence as observed in Agami (2023). Therefore, we propose new distance measures between persistence diagrams and persistence landscapes that are efficient in terms of time and also consider the time component inherent in time-series data.
Consider two time-series, ${x}=\{x_{t},~{}t=1,\ldots,T\}$ and ${y}=\{y_{t},~{}t=1,\ldots,T\}$ in $\mathbb{R}^{T}$. For a given length $\ell\in\{1,\dots,T\}$, we define the sequence of $m$ sub-series
$seq({x})$ $=$ $\{(x_{1},x_{2},\ldots,x_{\ell})_{1},(x_{1+\eta},x_{2+\eta},\ldots,x_{{\ell}+\eta})_{2},\ldots,(x_{T-{\ell}+1},x_{T-{\ell}+2},\ldots,x_{T})_{m}\},$
where $\eta\in\{1,\ldots,{\ell}\}$ represents the degree of overlapping (for $\eta=\ell$ the sub-series are non overlapping and $\eta=1$ represents maximum overlapping), and $T-{\ell}=\eta(m-1)$. We now define distance between ${x}$ and ${y}$ using their corresponding sub-series as follows,
Definition 3.1.
(Average $p$-Wasserstein distance ($AWD_{p}$)) For $1\leq p<\infty$, the average $p$-Wasserstein distance between the persistence diagrams ${\mathcal{D}}_{X}$ of ${x}$ and ${\mathcal{D}}_{Y}$ of ${y}$ is defined as follows
$AWD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{Y})=\sum_{i=1}^{m}w_{i}WD_{p}({\mathcal{D}}_{X_{i}},{\mathcal{D}}_{Y_{i}})$
where $WD_{p}({\mathcal{D}}_{X_{i}},{\mathcal{D}}_{Y_{i}})$ represent the $p$-Wasserstein distance between the persistence diagrams ${\mathcal{D}}_{X_{i}}$ and ${\mathcal{D}}_{Y_{i}}$ corresponding to the point clouds of $i$th components in $seq({x})$ and $seq({y})$. Here the weights $w_{i}$ are such that $w_{i}>0$ and $\sum_{i}w_{i}=1$. For instance, one can simply take $w_{i}=\frac{1}{m}$. For $p=\infty$, we also call the average Wasserstein distance as the average bottleneck distance ($ABD$).
Remark 3.1.
Computing $AWD_{p}$ is theoretically intuitive and computationally cheaper than calculating $WD_{p}$ between the two diagrams. Further, the choice of weights $w_{i}$ enables tuning of the influence of the preceding components of the time series when determining the distance.
Definition 3.2.
(Differenced $p$-Wasserstein distance ($DWD_{p}$)) The differenced $p$-Wasserstein distance ($DWD_{p}$) between the persistence diagrams ${\mathcal{D}}_{X}$ and ${\mathcal{D}}_{Y}$ of the point clouds ${X}$ and ${Y}$ of the time series ${x}$ and ${y}$ is defined as
$DWD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{Y})=WD_{p}({\mathcal{D}}_{X-Y},{\mathcal{D}}_{0}),$
where ${\mathcal{D}}_{X-Y}$ represents the persistence diagram corresponding to the point cloud obtained from the series ${x}-{y}$ and ${\mathcal{D}}_{0}$ is the diagram corresponding to a constant series (note that PDs are shift-invariant), containing only the diagonal Mileyko et al. (2011), so that $WD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{0})=\left(2^{-p}\sum_{a\in{\mathcal{D}}_{X}}(pers(a))^{p}\right)^{\frac{1}{p}}$, where $pers(a)$ denotes the persistence of $a=(b,d)\in{\mathcal{D}}_{x}$, given by $pers(a)=d-b$, with $b$ and $d$ denoting the birth and death time.
Clearly, for the toy example, $AWD$ and $DWD$ are non-zero. A similar distance measures could also be defined on the space of persistence landscapes. Let $\lambda_{x}$ and $\lambda_{y}$ denote the persistence landscapes for the time series $x$ and $y$, respectively. The average $p$-Landscape distance ($ALD_{p}$) and the differenced $p$-Landscape distance ($DLD_{p}$) between time series $x$ and $y$ as follows
[…]
where $\lambda_{x-y}$ is the persistence landscape corresponding to the point cloud of series $x-y$ and $\lambda_{x}^{i}$ and $\lambda_{y}^{i}$ are the persistence landscapes corresponding to the point clouds of $i$th components in $seq({x})$ and $seq({y})$, respectively.
3.1 Stability of Persistence diagrams under new distance measures
Having introduced new distance measures, it is essential to ensure that the mapping from data to a persistence diagram remains continuous as a data representation. Given that data is often noisy, it is crucial for the persistence diagrams to exhibit stability under perturbations in the data, that is, to demonstrate resilience to minor deformations in the data input. This section focuses on establishing the stability of persistence diagrams under the proposed distance measures.
Definition 3.3.
Let $X$ and $Y$ be two compact sets in ${\mathbb{R}}^{d}$ i.e., $X,Y\subset{\mathbb{R}}^{d}$. The Hausdorff distance $HD(X,Y)$ between $X$ and $Y$ is defined as
$HD(X,Y)=\inf\left\{\epsilon\geq 0\mid\text{$X\subseteq Y_{\epsilon}$ and $Y\subseteq X_{\epsilon}$}\right\},$ (8)
where $Y_{\epsilon}=\cup_{y\in Y}\{z\in{\mathbb{R}}^{d}\mid\|z-y\|\leq\epsilon\}$ denotes the generalized ball of radius $\epsilon$ around $Y$. Observe that
$X\subseteq Y_{\epsilon}$Longleftrightarrow for any $x\in X$ there exists some $y\in Y$ with $\|x-y\|\leq\epsilon$. (9)
The following theorem relates the Euclidean distance between two time series and Hausdorff distance between their corresponding point clouds.
Theorem 3.1.
Let ${x}$ and ${y}$ be two time series in ${\mathbb{R}}^{T}$. Let $X=\{x_{I_{1}},\dots,x_{I_{m}}\}$ and $Y=\{y_{I_{1}},\dots,y_{I_{m}}\}$ be their Takens’ embeddings in ${\mathbb{R}}^{d}$. Then
$HD(X,Y)\leq\|{x}-{y}\|_{2}.$ (10)
Proof.
For any $j=1,\dots,m$, we have $\|x_{I_{j}}-y_{I_{j}}\|^{2}_{2}=\sum_{t\in I_{j}}(x_{t}-y_{t})^{2}\leq\sum_{t=1}^{T}(x_{t}-y_{t})^{2}=\|{x}-{y}\|^{2}_{2}.$ In view of (9), we infer that $X\subseteq Y_{\|{x}-{y}\|}$ and $Y\subseteq X_{\|{x}-{y}\|}$. Combining this with (8), we arrive at Equation (10). ∎
We state the propositions required for our main theorem.
Proposition 3.2.
(Cohen-Steiner et al. (2007),Kusano et al. (2017),Chazal et al. (2014)) Let $(E,d_{E})$ be a metric space and $X$ and $Y$ be finite subsets of $E$. Then the persistence diagrams ${\mathcal{D}}_{X}$ and ${\mathcal{D}}_{Y}$ satisfy
$BD({\mathcal{D}}_{X},{\mathcal{D}}_{Y})\leq HD(X,Y).$
Proposition 3.3.
Cohen-Steiner et al. (2010),Kusano et al. (2017) Let $1\leq\hat{p}\leq p<\infty$, and ${\mathcal{D}}_{X}$ and ${\mathcal{D}}_{Y}$ be persistence diagrams whose degree $p$ total persistences [^7] Given a persistence diagram ${\mathcal{D}}_{X}=\{a_{j}=(b_{j},d_{j})\}$, the degree $p$ total persistence of ${\mathcal{D}}_{X}$ is defined as sum of persistence of all the points of degree $p$ in ${\mathcal{D}}_{X}$ i.e., $Pers_{p}({\mathcal{D}}_{X})=\sum_{j}(d_{j}-b_{j})$. are bounded from above. Then,
$WD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{Y})\leq\left(\frac{Pers_{\hat{p}}({\mathcal{D}}_{X})+Pers_{\hat{p}}({\mathcal{D}}_{Y})}{2}\right)^{\frac{1}{p}}BD({\mathcal{D}}_{X},{\mathcal{D}}_{Y})^{1-\frac{\hat{p}}{p}},$
where $Pers_{\hat{p}}({\mathcal{D}}_{X}):=\sum_{a\in{\mathcal{D}}_{X}}pers(x)^{\hat{p}}$ for $1\leq\hat{p}<\infty$.
Proposition 3.4.
Kusano et al. (2017) Let $E$ be a triangulable compact subspace in ${\mathbb{R}}^{d}$, ${X}$ be a finite subset of $E$, and $p>d$. Then,
$Pers_{p}({\mathcal{D}}_{X})\leq\frac{p}{p-d}C_{E}diam(E)^{p-d},$
where $C_{E}$ is a constant depending only on $E$.
Theorem 3.5.
Let ${x}$ and ${y}$ be two time series in ${\mathbb{R}}^{T}$. Let $X=\{x_{I_{1}},\dots,x_{I_{m}}\}$ and $Y=\{y_{I_{1}},\dots,y_{I_{m}}\}$ be their Takens’ embeddings in ${\mathbb{R}}^{d}$. Then
$ABD({\mathcal{D}}_{X},{\mathcal{D}}_{Y})\leq\|{x}-{y}\|_{2},$ (11)
$AWD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{Y})\leq\left(\frac{\hat{p}}{\hat{p}-d}C_{E}diam(E)^{\hat{p}-d}\right)^{\frac{1}{p}}\|{x}-{y}\|_{2}^{1-\frac{\hat{p}}{p}}.$ (12)
Proof.
By definition, for any choice of $\eta$ and $l$, $X_{i},~{}Y_{i};~{}i\in\{1,2,\ldots,m\}$, which represents the point cloud of $i$th component in $seq({x})$ and $seq({y})$, say ${x}_{i}$ and ${y}_{i}$, respectively, are again subsets of $E$. Therefore, using Proposition 3.2, we have
$BD({\mathcal{D}}_{X_{i}},{\mathcal{D}}_{Y_{i}})$$\leq$ $HD(X_{i},Y_{i})~{}~{}\forall~{}~{}i$
$\implies\sum_{i=1}^{m}w_{i}BD({\mathcal{D}}_{X_{i}},{\mathcal{D}}_{Y_{i}})$$\leq$ $\sum_{i=1}^{m}w_{i}HD(X_{i},Y_{i}),~{}~{}(\because~{}w_{i}>0,~{}\sum_{i=1}^{m}w_{i}=1)$
[…]
$\leq$ $\sum_{i=1}^{m}w_{i}\|{x}-{y}\|_{2},~{}~{}(\mbox{by definition of Euclidean distance})$
$\leq$ $\|{x}-{y}\|_{2}\sum_{i=1}^{m}w_{i}\leq\|{x}-{y}\|_{2}~{}~{}(\because\sum_{i=1}^{m}w_{i}=1.)$
Therefore, we have (11). Again, by definition, we have
$AWD_{p}({\mathcal{D}}_{X},{\mathcal{D}}_{Y})=\sum_{i=1}^{m}w_{i}WD_{p}({\mathcal{D}}_{X_{i}},{\mathcal{D}}_{Y_{i}})$
Using Proposition 3.3 and 3.4, we have
$\sum_{i=1}^{m}w_{i}WD_{p}({\mathcal{D}}_{X_{i}},{\mathcal{D}}_{Y_{i}})$$\leq$ $\sum_{i=1}^{m}w_{i}\left(\frac{Pers_{\hat{p}}({\mathcal{D}}_{X_{i}})+Pers_{\hat{p}}({\mathcal{D}}_{Y_{i}})}{2}\right)^{\frac{1}{p}}BD({\mathcal{D}}_{X_{i}},{\mathcal{D}}_{Y_{i}})^{1-\frac{\hat{p}}{p}}$
$\leq$ $\left(\frac{\hat{p}}{\hat{p}-d}C_{E}diam(E)^{\hat{p}-d}\right)^{\frac{1}{p}}\sum_{i=1}^{m}w_{i}BD({\mathcal{D}}_{X_{i}},{\mathcal{D}}_{Y_{i}})^{1-\frac{\hat{p}}{p}}$
$\leq$ $\left(\frac{\hat{p}}{\hat{p}-d}C_{E}diam(E)^{\hat{p}-d}\right)^{\frac{1}{p}}\sum_{i=1}^{m}w_{i}HD({X_{i}},{Y_{i}})^{1-\frac{\hat{p}}{p}}$
$\leq$ $\left(\frac{\hat{p}}{\hat{p}-d}C_{E}diam(E)^{\hat{p}-d}\right)^{\frac{1}{p}}\|{x}-{y}\|_{2}^{1-\frac{\hat{p}}{p}}.$
Therefore, we have result (12). ∎
Remark 3.2.
The $DWD_{p}$ is a special case of $WD_{p}$ with one component as zero $PD$. Therefore, the stability of persistence diagrams with respect to $DWD_{p}$ is trivial. Further, the stability of the proposed distances on persistence landscapes follows directly from the landscape stability theorems, Theorem 12 - Theorem 16 in Bubenik et al. (2015). Further, the distance metrics $AWD_{p},DWD_{p},ALD_{p}$ and $DLD_{p}$ are pseudo-metrics. The properties of non-negativity and symmetry are easily verified. Also, the triangular inequality property is satisfied since they are defined using the Wasserstein and $p$-landscape distance, which are again pseudo-metrics.
4 Our Methodology for Sparse Portfolios
Consider a stock-market index $I$ consisting of assets $i\in{\mathcal{I}}=\{1,2,\ldots,n\}$, in discrete time $t=0,1,2,\ldots,$, where $t$ represents the end of a trading day. At any time $t$, we denote the price of the index $I$ and any asset $i\in{\mathcal{I}}$ by $P_{0,t}$ and $P_{i,t}$ respectively. Further, we denote by $R_{i,t}$ and $r_{i,t}$, the simple-return and log-return of the asset $i$ over the period $[t-1,t]$ with $t=1,2,\ldots,$ and are obtained as follows,
$R_{i,t}=\frac{P_{i,t}-P_{i,t-1}}{P_{i,t-1}},~{}~{}r_{i,t}=\ln\,\frac{P_{i,t}}{P_{i,t-1}},\;\;i\in{\mathcal{I}}\cup\{0\};$
Now, consider an investor at time $T$, who desires to create a portfolio consisting of assets in ${\mathcal{I}}$, based on the historical observations, i.e., the time-series of prices (and hence of returns), $(P_{0,t},P_{1,t},\ldots,P_{n,t}),~{}t=0,1,2,\ldots,T$. We define a portfolio as an $n\times 1$ vector, denoted by $\bm{w}=(w_{1},w_{2},\ldots,w_{n})$, where $w_{i},i\in{\mathcal{I}}$ denotes the proportion of investment in the $i$th asset. We assume that the investor follows a so-called buy-and-hold strategy, i.e., she plans to hold this portfolio until the end of a specific investment horizon. Further, the set of all admissible portfolios is denoted by ${\mathcal{W}}$,
${\mathcal{W}}=\left\{{\bm{w}}=(w_{1},w_{2},\ldots,w_{n}):\sum_{i\in{\mathcal{I}}}w_{i}=1,w_{i}\geq 0~{}~{}\forall i\in{\mathcal{I}}\right\}.$
There are multiple ways to select the portfolio $\bm{w}$ depending on the goals and risk appetite of the investor. We discuss three of them, namely, mean-variance ($MV$) optimal portfolio, global minimum variance ($GMV$) portfolio, and index tracking ($IT$) portfolio with budget and short selling constraints.
Mean-Variance Portfolio In the standard Markowitz, i.e., the mean-variance portfolio optimization problem, the goal is to select a portfolio that delivers the highest expected returns for a given variance. Mathematically, the $MV$ portfolio corresponds to the solution of the following optimization problem,
$(MV)\qquad$ $\max_{\bm{w}\in{\mathcal{W}}}\begin{aligned} \bm{w}^{\prime}\bm{\mu}-\frac{\gamma}{2}\bm{w}^{\prime}\bm{\Sigma}\bm{w}\end{aligned}$ (13)
where $\bm{\mu}=(\mu_{1},\ldots,\mu_{n})$ is the sample mean vector and $\bm{\Sigma}$ is the sample variance-covariance matrix of the asset returns calculated over the observation period, $t=0,1,2\ldots,T$. Further, $\bm{w}^{\prime}\bm{\mu}$ and $\bm{w}^{\prime}\bm{\Sigma}\bm{w}$ denote the return and the variance of the portfolio. The parameter $\gamma$ is the risk aversion parameter that controls the trade-off between the portfolio’s risk and return. We follow DeMiguel et al. (2009b) and set $\gamma=1$ for the empirical analysis.
It is well documented in the literature that estimating mean vector $\bm{\mu}$ is more difficult than the variance-covariance matrix $\bm{\Sigma}$ of asset returns. Further, the impact of errors in $\bm{\mu}$ estimates is significantly larger than that of $\bm{\Sigma}$ on the portfolio $\bm{w}$. Therefore, we also consider a global minimum variance portfolio, which relies solely on estimates of $\Sigma$ and thus is less vulnerable to estimation error than $MV$ portfolio. The $GMV$ portfolio is the solution to the following optimization problem,
$(GV)\qquad$ $\min_{\bm{w}\in{\mathcal{W}}}\begin{aligned} \bm{w}^{\prime}\bm{\Sigma}\bm{w}\end{aligned}$ (14)
Index Tracking In an index tracking portfolio, the goal is to select the portfolio $\bm{w}$ that aims to mimic the performance of the index $I$ by minimizing the deviation of the portfolio return from the index return, the measure commonly known as tracking error $TE(\bm{w})$. The problem can be formulated as
$\min_{\bm{w}\in{\mathcal{W}}}\begin{aligned} TE({\bm{w}})\end{aligned}$ (15)
The most common choice for the tracking error $TE(\bm{w})$ in the literature is the mean of the squared deviations of the portfolio return and the index return, traditionally called the tracking error variance ($TEV$). We also adopt $TEV$ as the tracking error measure  [^8] One of the key benefits of employing this metric for tracking error is that the objective function is quadratic in the asset weights, allowing for using standard and very efficient quadratic programming algorithms to minimize it. A further advantage of this tracking error is that it avoids the issues associated with metrics based solely on the variance of the deviation since these measures ignore the possibility of a constant difference between the two returns (Beasley et al. (2003)). Instead of the squared deviations, the absolute error is also a popular choice in the literature (e.g., Prigent (2007); Rudolf et al. (1999)). This is advantageous because the optimization problem reduces to a linear program that can be efficiently solved. In practice, squared deviation is preferred over absolute errors as a statistical criterion, especially when large errors are undesirable. which is defined as
$\frac{1}{T}\sum_{t=1}^{T}(\sum_{i=1}^{n}w_{i}R_{i,t}-R_{0,t})^{2},$
where $R_{t}(\bm{w})=\sum_{i=1}^{n}w_{i}R_{i,t}$ denotes the portfolio return at time $t$.
4.1 TDA-based Sparse Portfolios
As discussed in the Introduction, various approaches for asset selection have been documented in the literature, especially for index-tracking and the Markowitz model to obtain sparse portfolios. In this paper, we adopt a strategy based on clustering that involves utilizing the distance measures introduced on persistence diagrams and landscapes. Building upon the new topology-based similarity measures, we seek to organize assets or clusters so that assets in the same cluster are similar and assets across the clusters are dissimilar in a topological sense. Once the clustering procedure is completed, we select the assets based on the following two strategies:
Strategy 1: This strategy corresponds to selecting assets that share the cluster with the index, intending to replicate the index’s performance. This choice is intuitive because the assets in the same cluster have a similar topological structure as quantified by the similarity matrix, thus narrowing down the number of assets potentially more suitable to mimic index performance. We then solve the tracking error optimization problem (15) on the selected assets to find the optimal weights.
Strategy 2: This strategy involves selecting assets for mean-variance and global minimum-variance portfolios. The underlying idea is to choose the representative assets from each of the clusters, driven by the intuition that this strategy promotes sparsity due to a smaller number of selected assets and diversity due to inter-cluster dissimilarity, as quantified by the similarity matrix. Once the assets are selected, we solve (13) and (14) to find the optimal weights. In this strategy, one needs to select the cluster representative. One possible way is to choose the representative based on criteria, such as the asset with the highest Sharpe ratio or the minimum standard deviation. However, this leads to a selection bias. Therefore, instead of imposing some random selection criteria, we propose selecting the cluster’s exemplar (representative center).
In this paper, we employ affinity propagation clustering (APC) Redmond et al. (2011) to obtain clusters. Our choice of adopting APC mainly stems from the requirement that the representative center of the clusters must be a part of the data set (refer to Strategy 2). Further, we should get the same result, i.e., the same clusters and centers, if we run our strategy on the same data set but on different platforms/software. Clustering techniques, such as $K$-Means and spectral clustering, assign random data points as initial cluster centers and shift these centers in each iteration. The drawback of these methods is that they yield different clustering sets upon each execution due to random initialization, and the centroids often lack a meaningful connection to actual variables in the dataset Jain et al. (1999). Affinity propagation addresses this issue by directly returning exemplars or centroids that correspond to actual data points.
Moreover, APC demonstrates superior accuracy compared to other clustering algorithms (Redmond et al. (2011)) such as K-means MacQueen et al. (1967), and spectral clustering Shi and Malik (2000); Murzin et al. (1995). Notably, APC doesn’t require the pre-specification of the number of clusters, unlike k-means and spectral clustering. Consequently, the similarity matrix is the only input needed for affinity propagation clustering. Finally, as we deal with time series of stock returns, which typically behaves like white noise, clustering algorithms such as K-means would not be suitable as demonstrated in Zheng et al. (2020).
Given the new topology-based distance measures, we define the following similarity matrices for the set of time series $x,y\in\mathbb{R}^{T}$ to be used as an input in affinity propagation clustering,
$K_{1}(x,y)$ $=e^{-\frac{AWD_{p}({\mathcal{D}}_{x},{\mathcal{D}}_{y})^{2}}{\sigma^{2}}},\quad$ $K_{2}(x,y)$ $=e^{-\frac{Wass_{p}({\mathcal{D}}_{x-y},{\mathcal{D}}_{0})^{2}}{\sigma^{2}}},$
$K_{3}(x,y)$ $=e^{-\frac{ALD_{p}(\lambda_{x},\lambda_{y})^{2}}{\sigma^{2}}},\quad$ $K_{4}(x,y)$ $=e^{-\frac{DLD_{p}(\lambda_{x},\lambda_{y})^{2}}{\sigma^{2}}},$ (16)
where $\sigma$ is a free parameter. [^9] These are kernels in strict sense Paulsen and Raghupathi (2016) and induce similarity matrices which are semi-positive definite and hence can be used as input for the clustering algorithms. It is worth highlighting that apart from the advantage of the new TDA-based distance measure over $p$-Wasserstein distance that it allows to compare time-series data as discussed in Section 3, it offers a substantial computational speedup as it reduces the computational complexity to compute the persistence landscape for a large time series by dividing it into computations over several sub-series. Further, this division into several sub-series accounts for heteroscedasticity and non-stationarity in the financial data.
For comparison purposes, we also utilize the following kernels based on traditional correlation measures, namely Spearman’s rank correlation and Pearson’s correlation,
[…]
where $d_{i}(x,y)=\sqrt{2(1-\rho_{i}(x,y))}~{}~{}i\in\{1,2\}$, with $\rho_{1}$ and $\rho_{2}$ denoting the Spearman’s rank correlation and Pearson’s correlation, respectively. [^10] The distance measure $d_{1}$ defined using Pearson’s correlation is the most commonly used distance measure in the literature Onnela et al. (2003). However, because these $x,y$ are logarithmic returns in our scenario, Spearman’s or Kendall’s rank correlation coefficient is a much better choice due to its robustness Zheng et al. (2020). As a result, we select both of these correlation measures for comparison. We use the same notation $K_{i}~{}i=1,\ldots,6$ to represent the kernels and the corresponding similarity matrix.
The kernels, $K_{1}$ to $K_{6}$, depend on hyper-parameter $\sigma$, which needs to be estimated from training data. This paper adopts a data-driven approach for estimating $\sigma$, specifically employing the local scaling algorithm Yang et al. (2008); Shang et al. (2012). More precisely, we use a local scaling parameter $\sigma_{ij}$ for each pair of input data $(x_{i},x_{j})$ defined as
$\sigma_{ij}=d(x_{i},x_{m}(x_{i}))d(x_{j},x_{m}(x_{j})),$
where $x_{m}(x_{\cdot})$ denotes the $m$th nearest neighbor of the input $x_{\cdot}$ in terms of distance measured by $d$ used in the corresponding kernel. [^11] The selection of $m$ is independent of scale. We take $m=7$ in our analysis since it provides good clustering results Yang et al. (2008). The adoption of diverse scaling parameters for each pair $(x_{i},x_{j})$ rather than a single scaling parameter $\sigma$ allows us to exploit the local statistics of the neighborhoods around the corresponding input pair, thereby facilitating a data-driven self-tuning of $\sigma$.
4.2 Performance Measures
Our goal is to evaluate the out-of-sample performance of the portfolio strategies mentioned above: Strategy 1 for index tracking and Strategy 2 for mean-variance and minimum variance optimal portfolios. We study the performance of these portfolios across various standard datasets in the literature on portfolio construction (more details on the datasets follow in the next section).
Given a daily time series of out-of-sample returns of optimal index tracking portfolio from strategy 1, we compute the five metrics. First, we calculate out-of-sample tracking error (TE): $\text{TE}={\frac{\sum_{t=1}^{T}(R_{t}(\bm{w})-R_{0,t})^{2}}{T}}.$
It measures how closely the optimally tracking portfolio replicates the benchmark index. Further, we also apply one-tailed $t$-test, with null hypothesis $H_{0}:\,\text{TE}_{s_{1}}-\text{TE}_{s_{2}}\leq 0$ against the alternative hypothesis $H_{a}:\text{TE}_{s_{1}}-\text{TE}_{s_{2}}>0$, to test whether the TE from strategy $s_{1}$ is statistically indistinguishable to that from strategy $s_{2}$.121212 The $t$-test is used to test if the mean $\mu$ of the population is ($\geq$  or  $\leq$) to some constant $\mu_{0}$ when the variance of the population is unknown. To test the following hypothesis $H_{0}:\mu\leq\mu_{0}~{}~{}~{}\mbox{against}~{}~{}~{}H_{a}:\mu>\mu_{0},$ the test statistics is given by $t=\frac{\bar{x}-\mu_{0}}{\frac{S}{\sqrt{N}}},$ where $S$ and $\bar{x}$ denote the sample variance and mean, respectively, and $N$ is the sample size. The critical region for the test is $t>t_{\alpha,N-1},$ where $\alpha$ is the significance level, and $t_{\alpha,N-1}$ denotes the critical value of the $t$-distribution with $N-1$ degrees of freedom. The $p$-value can be obtained as $P(t_{N-1}>t)$ and we reject the null hypothesis if $p-$value is less than $\alpha$. Second, we calculate excess mean return (EMR), defined as the mean of the difference between the out-of-sample returns from the optimal tracking portfolio and the returns of the benchmark index, that is, $\text{EMR}=\frac{\sum_{t=1}^{T}(R_{t}(\bm{w})-R_{0,t})}{T}.$ A positive (negative) EMR indicates that the tracking portfolio outperforms (underperforms) the benchmark index.
Third, to measure the extent to which the optimal tracking portfolio returns levels differ from that of the benchmark, we calculate Pearson’s correlation coefficient (COR) between the returns from the tracking portfolio and the benchmark index. A strategy for index tracking is appealing if its correlation is close to one. Fourth, we calculate the information ratio (IR), defined as the ratio of excess mean returns and tracking error, to measure the trade-off between excess mean return and tracking error: $\text{IR}=\frac{\text{EMR}}{\text{TE}}.$ Higher values of the information ratio are desirable.
Finally, to measure the tracking portfolio’s stability and understand the average wealth traded while holding (and re-balancing) the portfolio, we compute the portfolio turnover (TR). Formally, it is defined as the average absolute value of trades among $n$ stocks: $\text{TR}=\frac{1}{N}\sum_{r=1}^{N}\sum_{i=1}^{n}|w^{r}_{i}-w^{r-1}_{i}|,$ where $N$ is the total number of re-balancing windows, and $w^{r}_{i}$ and $w^{r-1}_{i}$ are respectively the weights of $i$th asset in $r$th and $(r-1)$th window. One can interpret $TR$ as a measure of the transaction costs for implementing the corresponding portfolio strategy. For instance, a lower TR means a lower transaction cost.
Similarly, given the time series of daily out-of-sample returns obtained from mean-variance and minimum variance optimal portfolios, we follow DeMiguel et al. (2009b) and compute four quantities to assess the performance of the portfolios. First, we measure the out-of-sample mean, which is given by $\mu={\sum_{t=1}^{T}\frac{R_{t}(\bm{w})}{T}}.$ To test whether the mean from the portfolio is statistically distinguishable, we also compute the p-value of the difference using the $t$-test. We apply a one-tailed $t$-test on null hypothesis $H_{0}:\,\mu_{s_{1}}-\mu_{s_{1}}=0$ against the alternative hypothesis $H_{a}:\mu_{s_{1}}-\mu_{s_{1}}>0$ to test if the out-of-sample mean return from $s_{1}$ is statistically significant from, $s_{2}$.
[…]
5 Empirical Results
This section presents our empirical analysis. Section 5.1 describes the data used in our study and introduces the rolling window scheme employed in the analysis. Sections 5.2 and 5.3 contain the performance results of our sparse index tracking and sparse mean-variance portfolios, respectively. Additionally, within these sections, we further validate the robustness of our findings through a performance analysis on the recent COVID-19 data.
5.1 Data
We consider the daily closing prices of constituents of the S&P 500 index, obtained from the Thomson Reuters datastream spanning a decade from 1st Dec 2009 to 31st Dec 2019 (2631 observations). We chose the data period according to the availability of a maximum number of constituents, resulting in 451 stocks in our dataset. We utilize daily log returns to obtain an input cloud for filtration and, consequently, to compute the distance measures for clustering. At the same time, we employ simple returns to solve the portfolio optimization problem.
For out-of-sample performance analysis, we follow a rolling window scheme. [^14] We look at the problem of constructing a sparse optimal portfolio for a period $1,\ldots,T$, which corresponds to the recent past (the in-sample period) and where the path of asset returns is known. The ultimate aim is to create portfolios close to optimum performance in the $T+1,\ldots,T+L$ period (the out-of-sample period). To elaborate, we begin our first window with a training period of 126 trading days (six months) as the in-sample period to create the optimal portfolio and track its performance over a testing period of 21 trading days (one month) without any re-balancing during this period. [^15] We follow the standard of 21 trading days in a month (see, for instance, Fastrich et al. (2014)). In the next window, we shift forward our training and testing period by one month to include more recent data and exclude the oldest data points such that a fixed size of the rolling window (seven months) is maintained. We then obtain an optimal portfolio on the new training period and track its performance over the new testing period. We repeat this procedure until the entire data sample is exhausted, consisting of 119 windows (test months) and thus resulting in a time series of portfolio returns, i.e., a one-month time series corresponding to each window. Finally, we analyze the performance of the optimal portfolios over the testing period, thereby offering a comprehensive assessment of the effectiveness of the filtering strategy. The rolling window procedure is also illustrated in Figure 5.
Figure 5:  Example of a rolling window scheme: In this illustration, the orange and blue hues indicate the in-sample and out-of-sample periods, respectively. For a fixed in-sample period of length $d_{1}$ and out-of-sample period of length $d_{2}$, the first in-sample period (rolling window 1) stands for $1,\ldots,d_{1}$ with corresponding out-of-sample period $d_{1}+1,\ldots,d_{1}+d_{2}$; the next in-sample period (rolling window 2) gets shifted by $d_{2}$ to obtain an in-sample window $1+d_{2}.\ldots,d_{1}+d_{2}$, with the accompanying out-of-sample period $d_{1}+d_{2}+1,\ldots,d_{1}+d_{2}+d_{2}$, and so on. For our analysis, $d_{1}=126$ and $d_{2}=21$.
To summarize, within a fixed rolling window, we obtain the time series of log returns for each constituent of the S&P 500 index during the in-sample period. We then follow the following steps:
  1. (i)
    Convert each time series into a point cloud to calculate kernels $K_{2}$ and $K_{4}$. Divide the time series into sub-series and convert each into a point cloud to calculate kernels $K_{1}$ and $K_{3}$.
  2. (ii)
    Use these point clouds to generate corresponding persistence diagrams and persistence landscapes.
  3. (iii)
    Calculate pair-wise distances between all possible pairs of stock return time-series, as defined in Section 3.
  4. (iv)
    Finally, obtain similarity matrices using the kernels $K_{i},~{}i=1,2,3,4$. The calculations for kernels $K_{5}$ and $K_{6}$ are straightforward.
  5. (v)
    With each of the obtained similarity matrices, apply affinity propagation clustering, resulting in clusters.
  6. (vi)
    Select assets from the obtained clusters using the proposed strategies, Strategy 1 and Strategy 2. Solve the respective optimization problems using simple returns of the constituents to get optimal portfolios.
  7. (vii)
    Finally, track the performance of these portfolios over the out-of-sample period.
5.2 Strategy 1: Sparse Index Tracking
In this subsection, we assess the performance of our TDA-based data-driven sparse index tracking portfolios. We denote the optimal tracking portfolio corresponding to a similarity matrix $K_{i}$ by $C_{i}$, $i=1,2,\ldots,6$. We also create the following portfolios from the literature as benchmarks for comparative analysis.
  • 1.
    IT with cardinality constraint: This portfolio corresponds to the optimization problem (15) with an additional constraint imposed on the maximum number of assets that can be selected within a tracking portfolio. Specifically, we impose the following restriction that the number of holdings with non-zero weights $\leq K_{max}$ where we set $K_{max}$ equal to the number of assets selected in the portfolio $C_{i},~{}i=1,2,\ldots,6$, to ensure a comparable number of asset for a fair comparison to the clustering based portfolios. Consequently, this yields six portfolios, denoted by $I_{i},~{}i=1,2,\ldots,6$. Given that the problem (15) with an additional constraint on the number of assets is a mixed-integer quadratic programming (MIQP) problem, it poses a challenging computational task. To tackle this computational bottleneck, we employ the Gurobi optimizer, which is recognized for its prowess in solving mixed-integer programming (MIP). For practical considerations, we follow Shu et al. (2020) and impose time limits of thirty minutes and one hour for solving each MIQP within every window. To distinguish, we denote the portfolios solved within a 30-minute time limit by $I_{i}^{(1)}$, while those solved with a one-hour time limit by $I_{i}^{(2)}$.
  • 2.
    IT with similarity based selection: Considering that the clustering-based portfolios $C_{i}$ are constructed by selecting assets from the clusters identified using similarity matrices, an alternative and more straightforward approach could be to eliminate the need for clustering altogether and choose the assets with the maximum similarity to the index. To settle this debate, we introduce an additional set of six portfolios, denoted by $MS_{i},i=1,2,\ldots,6$. Each $MS_{i}$ corresponds to the portfolio obtained by solving (15) on top $M$ assets selected based on maximum similarity with the index, as measured by the similarity matrix $K_{i}$. We consider the maximum similarity criteria with $M=20$ as a benchmark.
  • 3.
[…]
    $K_{7}(X,Y)=-||X-Y||_{2}^{2}.$
    We denote this portfolio by $C_{7}$.
Before analyzing the performance of the portfolios described above, we first examine the number of assets in the cluster containing the index resulting from APC based on the seven similarity matrices, $K_{i},~{}i=1,2,\ldots,7$, following our strategy 1. The mean and median number of assets in all windows are reported in Table 1. There are several observations. The number of clusters formed with TDA-based similarity matrices $K_{1}$ to $K_{4}$, as well as the correlation-based similarity matrices $K_{5}$ and $K_{6}$, is consistent over the sample period. A small number, on average, of assets in these clusters results in sparse portfolios as desired. For instance, the similarity matrix $K_{2}$ selects 40 assets (approximately 8%) on average, while the correlation-based matrix $K_{5}$ (or $K_{6}$) selects a further smaller number of assets. In contrast, the similarity matrix $K_{7}$, on average, selects almost 57% of assets, which is not desirable, and consequently, we drop the portfolio $C_{7}$ from our analysis. [^17] In Appendix 6.2, we present the study of the properties of the selected assets with respect to the benchmark index.
Table 1: Mean and Median number of assets sharing cluster with the index
Similarity matrix $K_{1}$ $K_{2}$ $K_{3}$ $K_{4}$ $K_{5}$ $K_{6}$ $K_{7}$
Mean 7 40 6 12 18 18 255
Median 6 38 5 12 18 17 256
Table 2: The out-of-sample performance of clustering based portfolios $C_{1}-C_{6}$, maximum similarity based portfolios $MS_{1}-MS_{6}$ and portfolios from TEV model with cardinality constraint $I_{1t_{1}}-I_{6t_{1}}$ and $I_{1t_{2}}-I_{6t_{2}}$.
$C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ $C_{5}$ $C_{6}$
TE 2.62E-05 5.14E-06 2.79E-05 5.00E-06 1.21E-05 1.08E-05
EMR 1.44E-04 1.30E-04 -1.05E-05 3.40E-05 9.93E-05 9.03E-05
COR 8.36E-01 9.68E-01 8.32E-01 9.70E-01 9.39E-01 9.44E-01
INFO 2.81E-02 5.71E-02 -1.98E-03 1.52E-02 2.86E-02 2.75E-02
TR 1.47 1.10 1.70 1.58 1.18 1.08
$MS_{1}$ $MS_{2}$ $MS_{3}$ $MS_{4}$ $MS_{5}$ $MS_{6}$
TE 1.01E-05 7.61E-06 1.18E-05 6.85E-06 8.79E-06 8.76E-06
EMR 3.71E-05 5.74E-05 -3.90E-05 6.62E-05 6.00E-05 5.91E-05
COR 9.04E-01 9.36E-01 8.85E-01 9.37E-01 9.37E-01 9.41E-01
INFO 9.55E-03 1.77E-02 -9.07E-03 2.02E-02 1.63E-02 1.71E-02
TR 1.02 7.74E-01 1.34 1.47 9.54E-01 8.50E-01
$I_{1t_{1}}$ $I_{2t_{1}}$ $I_{3t_{1}}$ $I_{4t_{1}}$ $I_{5t_{1}}$ $I_{6t_{1}}$
TE 5.70E-05 3.14E-05 3.20E-05 5.41E-05 2.55E-05 4.67E-05
EMR -5.81E-05 1.90E-04 -1.54E-04 2.25E-04 -3.14E-05 -1.67E-05
COR 7.57E-01 8.52E-01 8.43E-01 7.66E-01 8.75E-01 8.01E-01
INFO -7.69E-03 3.39E-02 -2.72E-02 3.06E-02 -6.22E-03 -2.45E-03
TR 1.72 1.61 1.65 1.64 1.68 1.62
$I_{1t_{2}}$ $I_{2t_{2}}$ $I_{3t_{2}}$ $I_{4t_{2}}$ $I_{5t_{2}}$ $I_{6t_{2}}$
TE 1.34E-05 2.45E-06 1.28E-05 8.00E-06 6.57E-06 5.13E-06
EMR 1.36E-04 3.11E-05 -3.70E-06 9.91E-06 7.48E-05 4.76E-05
COR 9.26E-01 9.85E-01 9.32E-01 9.54E-01 9.62E-01 9.71E-01
INFO 3.73E-02 1.99E-02 -1.03E-03 3.50E-03 2.92E-02 2.10E-02
TR 1.63 1.58 1.66 1.71 1.68 1.66
Table 2 presents the out-of-sample performance of all portfolios in terms of five performance measures, i.e., correlation with the index (COR), tracking error (TE), excess mean return (EMR), information ratio (INFO), and the turnover ratio (TR). The best and second-best values among models are highlighted in bold and italics, respectively. Given that it is challenging to compare all the models simultaneously, we first identify the best-performing portfolio among those constructed based on clustering using TDA-based distance measures. Subsequently, we comprehensively compare these selected portfolios to those based on maximum similarity criteria and cardinality constraints.
First, it is evident that all but $C_{3}$ portfolios generate positive EMR. Although an ideal tracking portfolio should have zero excess mean return, it is rarely observed in practice. We usually observe both positive and negative values with a positive (negative) EMR indicating that the tracking portfolio outperforms (under-performs) the benchmark index. A positive return over the benchmark is desirable as it can certainly offset the negative impact of transaction costs. Second, in terms of the tracking performance, the portfolios $C_{4}$ and $C_{2}$ outperform the other TDA-based clustering portfolios in terms of out-of-sample tracking error and result in the highest correlation of 96.8% and 97%, respectively, with the index.
Furthermore, the tracking error from $C_{5}$ and $C_{6}$ is relatively high compared to the tracking error from $C_{2}$ and $C_{4}$, demonstrating the proposed kernels’ efficacy in filtering out more identical assets than the conventional kernels. Third, $C_{2}$ excels $C_{4}$ as it attains the highest information ratio and maintains a trade-off between the return over and above the benchmark index and the tracking error. Last but not least, we compare the turnover values that reflect the potential transaction costs associated with portfolio re-balancing, with higher values showing more re-balancing, hence higher transaction costs, and vice versa. The turnover values show that the investor needs to re-balance fewer positions for $C_{6}$ and $C_{2}$ compared to the re-balancing for $C_{4}$ portfolios from one investment period to the other. Finally, we conclude that both $C_{2}$ and $C_{4}$ outperform the other clustering-based portfolios, with $C_{2}$ outperforming $C_{4}$ in terms of turnover ratio and information ratio.
Next, we observe that both the portfolios $C_{2}$ and $C_{4}$ outperform the portfolios based on maximum similarity selection criteria with $M=10$ in terms of the tracking error and correlation with the index. Further, $C_{2}$ also beats the maximum similarity portfolios in terms of EMR, while $C_{4}$ fails to do so. One can attribute the smaller values of TR for maximum similarity portfolio compared to $C_{2}$ and $C_{4}$ to the fact the $MS$ portfolios consist of only ten assets and hence the less re-balancing.
Comparing the performance of the portfolios $I_{1t_{1}}-I_{6t_{1}}$ to that of the portfolios $I_{1t_{2}}-I_{6t_{2}}$, we find that the latter set of portfolios outperforms the former in terms of TE, EMR, and correlation with the index. In contrast, for other performance measures, INFO and TR, there is no clear domination. Further, the portfolios $I_{1t_{1}}-I_{6t_{1}}$ perform poor than the TDA-based sparse portfolios. When we compare the clustering-based portfolios to portfolios represented as $I_{1t_{2}}-I_{6t_{2}}$, we observe that the tracking error from $C_{2}$ is greater than that of $I_{2t_{2}}$, but the turnover is relatively low. This higher tracking error in the case of $C_{2}$ is quite intuitive. We minimize tracking error by utilizing all available assets when considering $I_{2t_{2}}$. In contrast, with $C_{2}$, we first filter out a subset of assets and then minimize the tracking error on this reduced set, resulting in a sub-optimal solution. However, it’s important to note that the former approach incurs higher transaction costs because, in each evaluation window, the optimization problem is solved anew using all assets, leading to frequent portfolio re-balancing.
On the other hand, for $C_{2}$, the initial filtering step imparts stability to the problem, resulting in less frequent portfolio re-balancing and, consequently, lower transaction costs. In conclusion, no consensus exists for the best among $C_{2}$, $C_{4}$, and $I_{1t_{2}}-I_{6t_{2}}$. There is a trade-off between the tracking error and the transaction costs. Nonetheless, clustering-based portfolios require less computational time when compared to cardinality-constrained portfolios. To find the best portfolio, we further analyze their performance. That is, given the superior performance of the portfolios $C_{2}$, $C_{4}$ and $I_{1t_{2}}-I_{6t_{2}}$, we now compute the mean ratio of the in-sample tracking error and out-of-sample tracking error  [^18] We calculate the ratio of in-sample tracking error and out-of-sample tracking error in each window and then take average to find the mean Canakgoz and Beasley (2009). We also compare it to the mean ratio of the portfolio $TE_{all}$, which corresponds to a complete replication index tracking portfolio, i.e., a tracking portfolio constructed using all the assets constituting the index without penalty terms. A comparison of the mean ratios allows us to investigate the severity of the estimation error. Table 3 shows the average value of in-sample and out-of-sample tracking error across the windows and the average value of the corresponding ratios. A ratio of less than one across all portfolios indicates that, on average, there is a performance decline when transitioning from the in-sample period to the out-of-sample period. It is not surprising since the portfolio is optimized over the in-sample period, thus generating a smaller tracking error, while the same portfolio composition, when tracked over the out-of-sample period, the data that is unseen before. Nevertheless, it is remarkable that performance degradation is not excessively high, particularly for portfolios based on clustering techniques. This reaffirms the robustness of clustering-based portfolios in both the in-sample and out-of-sample periods.
Table 3: In-sample and out-of-sample robustness check. In-TE is the average in-sample value of the tracking error across the windows, Out-TE is the average out-of-sample value of the tracking error across the windows and Mean ratio is the mean (taken across the windows) value of ratio of in-sample and out-of-sample tracking error. $C_{2}$ and $C_{4}$ are clustering based portfolios, $I_{1t_{2}}-I_{6t_{2}}$ are portfolios from cardinality constraint model and $TE_{all}$ is the portfolio build using all the assets.
Clustering Cardinality constraint All
$C_{2}$ $C_{4}$ $I_{1t_{2}}$ $I_{2t_{2}}$ $I_{3t_{2}}$ $I_{4t_{2}}$ $I_{5t_{2}}$ $I_{6t_{2}}$ $TE_{all}$
In TE 2.4E-06 2.2E-06 5.1E-06 1.7E-07 4.8E-06 2.1E-06 2.4E-06 1.0E-06 5.4E-15
Out TE 5.1E-06 5.0E-06 1.3E-05 2.4E-06 1.3E-05 8.0E-06 6.6E-06 5.1E-06 9.0E-07
Mean ratio 5.5E-01 5.2E-01 4.4E-01 7.4E-02 4.7E-01 3.0E-01 2.3E-01 2.3E-01 7.1E-09
Remark 5.1.
To further validate the exceptional performance of our top-performing TDA-based sparse portfolio, $C_{2}$, we compare it to the tracking portfolios constructed using the co-integration technique, a methodology known for its effectiveness in index-tracking problems Alexander and Dimitriu (2005), Sant’Anna et al. (2017). We denote this portfolio by $CO$. The idea underlying the co-integration index tracking portfolio revolves around determining the optimal weights for a given set of assets $S_{1},\ldots,S_{n}$ that achieve the highest level of co-integration with the benchmark. Nevertheless, selecting the parameter $n$ is vital and demands careful consideration. For instance, to obtain a stable co-integrating vector of at least $n$ stocks out of $N$ stocks in the index, a mechanical search would entail examining an extensive array of portfolios - specifically, $N!/(N-n)!n!$ portfolios - as potential candidates for backtesting. However, in our paper, we adopt the approach presented in Sant’Anna et al. (2017) for constructing a co-integration-based index-tracking portfolio. More specifically, we start by randomly selecting 40 stocks, a choice motivated by our intention to compare it to the $C_{2}$ portfolio, which, on average, consists of 40 stocks. We run a regression analysis (for a comprehensive understanding of the co-integration model, please refer to Sant’Anna et al. (2017)) on the in-sample data and select this portfolio as a candidate if the residuals from the regression are stationary, as indicated by the Augmented Dickey-Fuller (ADF) test. The stationary residuals would ascertain the presence of co-integration between the index and asset price series. Otherwise, we discard the portfolio under consideration. We repeat this process for 100,000 regressions conducted sequentially for each portfolio. Each regression employs a unique combination of 40 assets. The final portfolio chosen is determined by the regression that yields the smallest sum of squared residuals. Our findings suggest that the $C_{2}$ portfolio outperforms the $CO$ portfolio in terms of all the performance metrics as reported in Table 4. For instance, the $C_{2}$ portfolio achieves a significantly lower tracking error of 5.14E-06 in contrast to 1.57E-05 achieved by the $CO$ portfolio. Similarly, the correlation of 0.968 with the index for the $C_{2}$ portfolio is significantly higher than 0.924 for the $CO$ portfolio. These findings suggest that TDA-based sparse index tracking portfolios perform better than traditional portfolios based on correlation and co-integration with the index.
Table 4: The out-of-sample performance of $C_{2}$ portfolio constructed using Strategy 1 with kernel 2 and the $CO$ portfolio obtained from co-integration based index tracking model.
$C_{2}$ $CO$
TE 5.14E-06 1.57E-05
EMR 1.30E-04 5.06E-05
COR 9.68E-01 9.24E-01
INFO 5.71E-02 1.28E-02
TR 1.10E+00 2.38E+00
Overall, we observed an improvement in the index tracking problem using the TDA-based kernels with strategy 1. The key concern, though, is the robustness of the strategy. The market fluctuations are not deterministic and cannot be captured. So, we analyze the performance of TDA-based kernels during the crisis period. In particular, we look at the recent Covid-19 pandemic.
Covid-19 robustness check: As a robustness check, Table 5 shows the out-of-sample performance of tracking portfolios for data from the recent COVID-19 pandemic, July 2019 to mid-November 2020. Again, we use the daily prices of the S&P500 index and its 502 constituents as available in this period. We adopt the rolling window strategy with 21 days (approx one month) in-sample period and five days (approx one week) out-of-sample period. Among TDA-based portfolios, we observe that $C_{2}$ and $C_{4}$ still have the lowest and statistically significant tracking errors. Also, $C_{2}$ attains minimum turnover compared to other portfolios.
In contrast to the observations in the stable period, the maximum similarity-based portfolio $MS_{2}$ outperforms the best TDA-based portfolios, i.e., $C_{2}$ and $C_{4}$. For instance, $MS_{2}$ achieves a tracking error of 3.05E-05, better than 3.86E-05 and 3.53E-05 obtained by $C_{2}$ and $C_{4}$. Similarly, the turnover ratio 1.34 by $MS_{2}$ outperforms the ratio obtained by $C_{2}$ and $C_{4}$ of 1.66 and 1.94, respectively. The comparison of TDA-based portfolios against cardinality-based portfolios follows a pattern similar to the findings in the stable period in that there is no clear relation between them. However, considering the trade-off between the performance and the computational time, clustering-based portfolios outperform.
Table 5: The out-of-sample performance of clustering based portfolios $C_{1}-C_{6}$, maximum similarity based portfolios $MS_{1}-MS_{6}$ and portfolios from TEV model with cardinality constraint $I_{1t_{1}}-I_{6t_{1}}$ and $I_{1t_{2}}-I_{6t_{2}}$ on the Covid-19 pandemic period.
$C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ $C_{5}$ $C_{6}$
TE 7.88E-05 3.86E-05 8.96E-05 3.53E-05 8.59E-05 1.24E-04
EMR -5.85E-05 -1.51E-04 -5.68E-04 -2.44E-04 -5.36E-04 9.48E-04
COR 9.09E-01 9.52E-01 8.88E-01 9.53E-01 8.91E-01 8.58E-01
INFO -6.59E-03 -2.43E-02 -6.00E-02 -4.10E-02 -5.78E-02 8.53E-02
TR 1.82 1.66 1.83 1.94 1.68 1.66
$MS_{1}$ $MS_{2}$ $MS_{3}$ $MS_{4}$ $MS_{5}$ $MS_{6}$
TE 7.02E-05 3.05E-05 1.06E-04 6.97E-05 6.34E-05 4.56E-05
EMR -7.98E-04 1.47E-04 1.17E-04 -6.86E-06 3.91E-04 2.78E-04
COR 9.34E-01 9.61E-01 8.62E-01 9.21E-01 9.26E-01 9.49E-01
INFO -9.52E-02 2.66E-02 1.13E-02 -8.22E-04 4.90E-02 4.12E-02
TR 1.74 1.34 1.84 1.72 1.44 1.42
$I_{1t_{1}}$ $I_{2t_{1}}$ $I_{3t_{1}}$ $I_{4t_{1}}$ $I_{5t_{1}}$ $I_{6t_{1}}$
TE 5.93E-05 2.90E-05 3.91E-05 3.27E-05 3.56E-05 4.67E-05
EMR 4.29E-04 -1.63E-04 3.24E-04 -2.49E-04 4.97E-04 3.82E-04
COR 9.29E-01 9.62E-01 9.52E-01 9.58E-01 9.55E-01 9.45E-01
INFO 5.57E-02 -3.03E-02 5.18E-02 -4.35E-02 8.33E-02 5.58E-02
TR 1.84 1.92 1.84 1.83 1.93 1.91
$I_{1t_{2}}$ $I_{2t_{2}}$ $I_{3t_{2}}$ $I_{4t_{2}}$ $I_{5t_{2}}$ $I_{6t_{2}}$
TE 5.80E-05 2.91E-05 3.99E-05 3.27E-05 3.58E-05 3.34E-05
EMR 3.19E-04 -1.71E-04 4.09E-04 -2.49E-04 5.54E-04 4.73E-04
COR 9.29E-01 9.62E-01 9.51E-01 9.58E-01 9.55E-01 9.55E-01
INFO 4.18E-02 -3.17E-02 6.48E-02 -4.35E-02 9.27E-02 8.18E-02
TR 1.829 1.918 1.827 1.830 1.920 1.917
Case Study (Publicly available data for Index tracking model): For completeness, we also test the performance of our TDA-based sparse index tracking portfolios on the benchmark data sets from Beasley et al. (2003); Canakgoz and Beasley (2009) for index tracking problems. These data sets are drawn from several major world equity indices, and thus, analyzing the performance of these data sets gives additional evidence on the adequacy and effectiveness of the TDA-based portfolios.
Table 6: Indices used for empirical analysis
S. No. Index Country Total Number Constituents
of constituents considered
1. FTSE 100 UK 100 89
2. Nikkei 225 Japan 225 225
3. DAX 100 Germany 100 85
4. Hang-Seng Hong Kong 50 31
5. S&P 100 USA 100 98
6. S&P 500 USA 500 457
7. Russell 2000 USA 2000 1318
8. Russell 3000 USA 3000 2151
Our data sets belong to the OR-Library and are publicly available at http://people.brunel.ac.uk/~mastjjb/jeb/info.html. There are currently eight benchmark instances. Each instance corresponds to the securities composing a specific market index. Specifically, the following market indices are considered: Hang Seng (Hong Kong), DAX100 (Germany), FTSE100 (United Kingdom), S&P100 (USA), Nikkei225 (Japan), S&P500 (USA), Russell2000 (USA) and Russell3000 (USA). The data set comprises 291 weekly closing prices for each of the indices spanning March 1992 to September 1997. Table 6 gives the list of indices, their actual number of constituents, and the number of constituents considered for this study.
We follow Beasley et al. (2003) and consider the first 145 weekly prices as in-sample observations and the following 146 weeks as out-of-sample periods to evaluate the performance of the models. We do not require a rolling window approach here, as there is only one training and testing period. Consequently, instead of reporting transaction costs, we find the Hinderfindahl index, a diversity index calculated as $\text{HHI}=\sum_{i=1}^{n}w_{i}^{2}.$ In the case of a naïve portfolio (which offers maximum diversification), the minimum value of HHI is obtained and is equal to $\frac{1}{n}$. Smaller values of HHI are desirable.
Table 7: Out-of-sample performance of strategy 1 on data set from Beasley OR Library. Here $M$ denotes the number of assets filtered as clustering result, i.e., the number of assets sharing the same cluster with the index
T1 $C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ $C_{5}$ $C_{6}$ $TE_{all}$ T5 $C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ $C_{5}$ $C_{6}$ $TE_{all}$
EMR -4.61E-05 9.65E-05 -4.85E-04 -3.76E-05 2.59E-04 9.16E-04 9.96E-05 EMR -1.32E-03 -7.80E-04 -3.65E-05 -1.04E-03 -9.76E-04 -1.24E-03 -3.68E-04
TE 1.27E-05 1.46E-05 1.84E-04 9.82E-05 1.96E-05 4.83E-05 7.30E-06 TE 5.85E-05 4.58E-05 1.53E-04 8.80E-05 6.34E-05 6.91E-05 2.96E-05
COR 9.92E-01 9.91E-01 9.27E-01 9.36E-01 9.89E-01 9.79E-01 9.95E-01 COR 9.66E-01 9.70E-01 8.94E-01 9.43E-01 9.71E-01 9.61E-01 9.84E-01
HHI 8.81E-02 8.82E-02 1.96E-01 1.56E-01 9.57E-02 1.21E-01 7.05E-02 HHI 2.90E-02 3.85E-02 2.60E-01 9.01E-02 4.92E-02 4.80E-02 1.37E-02
M 19 18 6 8 18 15 31 M 62 54 5 15 57 46 225
T2 T6
EMR 3.32E-04 1.39E-04 -2.07E-04 7.67E-04 3.43E-04 1.78E-04 -2.66E-05 EMR 3.08E-03 2.15E-03 3.55E-03 2.34E-03 2.32E-03 1.97E-03 2.22E-03
TE 1.12E-04 7.96E-05 1.79E-04 1.14E-04 9.76E-05 8.77E-05 5.76E-05 TE 7.64E-04 2.85E-04 7.30E-04 3.24E-04 5.78E-04 4.33E-04 1.03E-04
COR 8.61E-01 9.06E-01 8.05E-01 8.67E-01 8.86E-01 8.95E-01 9.30E-01 COR 6.37E-01 7.89E-01 4.86E-01 7.71E-01 7.50E-01 8.01E-01 9.28E-01
HHI 1.20E-01 1.12E-01 3.42E-01 1.58E-01 1.55E-01 1.26E-01 3.87E-02 HHI 3.11E-01 5.67E-02 2.51E-01 9.50E-02 2.65E-01 2.14E-01 1.80E-02
M 27 12 4 8 8 10 85 M 4 40 7 24 6 7 457
T3 T7
EMR -8.44E-04 -9.27E-04 2.47E-04 -9.36E-04 -1.23E-03 -1.58E-03 -1.72E-04 EMR 2.32E-03 2.24E-03 3.02E-03 2.50E-04 6.02E-04 -3.54E-04 1.79E-03
TE 7.94E-05 2.95E-05 8.27E-05 5.12E-05 5.25E-05 4.16E-05 7.83E-06 TE 5.65E-04 3.96E-04 7.80E-04 6.66E-04 5.39E-04 6.09E-04 2.14E-04
COR 8.38E-01 9.36E-01 8.24E-01 8.91E-01 8.84E-01 9.11E-01 9.83E-01 COR 7.07E-01 7.74E-01 5.20E-01 5.65E-01 7.18E-01 6.89E-01 8.93E-01
HHI 1.05E-01 4.36E-02 2.31E-01 9.42E-02 5.85E-02 6.43E-02 2.26E-02 HHI 1.09E-01 3.54E-02 1.53E-01 4.88E-02 6.48E-02 9.56E-02 1.33E-02
M 11 35 5 14 25 24 89 M 14 63 8 47 40 20 1318
T4 T8
EMR -4.47E-04 -2.75E-04 -9.72E-04 -5.62E-04 -4.80E-04 -7.23E-04 -1.26E-04 EMR 1.43E-03 1.71E-03 3.69E-03 1.76E-03 1.67E-03 2.23E-03 2.52E-03
TE 3.63E-04 3.86E-05 1.65E-04 5.31E-05 6.69E-05 7.03E-05 9.35E-06 TE 6.49E-04 2.33E-04 7.31E-04 3.18E-04 5.93E-04 5.26E-04 1.14E-04
COR 6.30E-01 9.35E-01 7.44E-01 9.16E-01 8.93E-01 8.95E-01 9.85E-01 COR 6.09E-01 8.32E-01 3.21E-01 7.70E-01 7.17E-01 7.38E-01 9.26E-01
HHI 3.69E-01 7.24E-02 2.21E-01 9.87E-02 1.06E-01 1.19E-01 2.70E-02 HHI 1.49E-01 4.40E-02 2.09E-01 5.49E-02 2.29E-01 1.62E-01 1.63E-02
M 5 24 5 14 15 13 98 M 8 62 8 50 8 13 2151
Table 7 reports the performance of TDA-based sparse index tracking portfolios $C_{1}$ to $C_{6}$ on the eight indices and the fully-replicating index tracking portfolio, $TE_{all}$. There are several findings. First, our $C_{2}$ portfolio achieves the lowest tracking error when contrasted against the other TDA-based portfolios for all the indices under consideration. What’s more, it is striking that the tracking error stemming from $C_{2}$ is statistically insignificant compared to the tracking error emanating from the $TE_{all}$ portfolio for five of the eight indices. Second, the correlation between the $C_{2}$ portfolios and the respective market index is notably higher than that observed in other TDA-based portfolios for five of the eight indices. Even for the index where the correlation from the $C_{2}$ portfolios is not the highest, it remains remarkably close to the highest correlation value. Third, although the portfolios generate a negative excess mean return for some of the indices, it is interesting that even in these instances, the $C_{2}$ portfolios consistently outperform their counterparts, namely the $C_{5}$ and $C_{6}$ portfolios, which are constructed using conventional similarity matrices. Fourth, it is remarkable that the $C_{2}$ portfolios are well-diversified, as indicated by the Herfindahl-Hirschman Index (HHI), with $C_{2}$ portfolios achieving the lowest values among the portfolios. It is essential to underscore that the lower values of HHI for $C_{2}$ attributes to the specific assets chosen based on the clustering with a similarity measure based on the kernel $K_{2}$ and is not merely a by-product of the higher value of $M$ (number of assets selected) as evidenced by the HHI values for T5. These findings indicate that the based sparse index tracking portfolio, $C_{2}$, consistently outperforms the other index tracking portfolios under consideration.
Remark 5.2.
We also compare the performance of our best-performing TDA-based sparse portfolio, $C_{2}$, to the index tracking portfolio proposed in Beasley et al. (2003), which we denote by $C_{beas}$. It is worth noting that our TDA-based sparse portfolio, unlike $C_{beas}$, doesn’t require the number of assets to be included in the portfolio. Since the composition of the optimal index tracking portfolios in Beasley et al. (2003) is not publicly available, we directly use the tracking error values reported in that paper. Further, the tracking error used in Beasley et al. (2003) is
$TE_{beas}=\frac{1}{T}\sqrt{\sum_{j=1}^{T}(R_{j}(\bm{w})-R_{0,j})^{2}}.$
Therefore, for a fair comparison, we also compute the tracking error, $TE_{beas}$, for our portfolio, $C_{2}$. Table 8 presents the out-of-sample $TE_{beas}$ for $C_{2}$ and $C_{beas}$ for data sets T1-T5, the indices used in Beasley et al. (2003). Interestingly, our $C_{2}$ portfolio outperforms $C_{beas}$ for all except T4 indices in terms of the tracking error $T_{beas}$.
Table 8: Out-of-sample value of tracking error $TE_{beas}$ for $C_{2}$ portfolio and $C_{beas}$ portfolio from $Beasley_{IT}$ model. Best value is highlighted in Bold
Data $C_{2}$ $C_{beas}$
T1 2.96E-04 1.27E-03
T2 8.77E-04 2.05E-03
T3 7.40E-04 9.58E-04
T4 1.58E-03 1.03E-03
T5 6.35E-04 8.21E-04
The empirical study demonstrates the efficacy of the TDA-based kernel $K_{2}$ in facilitating the superior asset selection for the index-tracking model. Further, the kernel $K_{2}$, when integrated with affinity propagation clustering, results in an entirely data-driven approach, thus eliminating the need to pre-specify the number of assets to be included in the tracking portfolio. Therefore, our proposed TDA based strategy holds promise in enhancing asset allocation decisions, particularly in the context of index tracking.
5.3 Strategy 2: Sparse Mean-variance and Global Minimum Variance Portfolios
Next, we assess the performance of our TDA-based data-driven mean-variance and global minimum variance optimal portfolios. More precisely, for a fixed window, we follow Strategy 2, i.e., given the similarity matrices $K_{1}$ to $K_{6}$, we perform affinity propagation clustering and select the exemplars, then construct the optimal portfolios solving problems (13) and (14) on the filtered stocks. We denote the optimal mean-variance and the global minimum variance portfolios corresponding to a similarity matrix $K_{i}$ by $MV_{i}$ and $GMV_{i}$, respectively, where $i=1,2,\ldots,6$. We also create the following portfolios from the literature as benchmarks for comparative analysis.
  • 1.
    MV and GMV with all assets: These portfolios correspond to the optimal mean-variance and global minimum variance portfolios created by solving (13) and (14) with all of the assets that make up the benchmark index (without clustering). We denote these portfolios by $MV_{A}$ and $GMV_{A}$, respectively.
  • 2.
    Naive portfolio: This portfolio corresponds to the equal-weighted portfolio.
Before analyzing the performance of optimal portfolios, we first observe that the median number of assets (exemplars) for similarity matrices $K1-K6$ are 60, 64, 74, 86, 73, and 76, thus confirming that the resulting portfolios are indeed sparse.
Panel A in Table 9 shows the out-of-sample performance of the TDA-based sparse mean-variance portfolios. We observe that the portfolio $MV2$ achieves the maximum values for mean, SR, and CEQ, and the minimum value of TR, thus outperforming the other portfolios, including TDA-based portfolios and the benchmark portfolios, namely, $MVA$, the naïve strategy and the benchmark index. For instance, $MV2$ achieves the daily Sharpe ratio of .0733 (.3359 per month / 1.1636 per annum) compared to 0.688 and 0.521 achieved by $MVA$ and $SP500$. Further, the least value of the turnover ratio for $MV2$ indicates less re-balancing across the windows and the formation of stable portfolios. In contrast to the performance of TDA-based index-tracking portfolios, $MV4$ fails to consistently beat other TDA-based portfolios as well as the benchmark portfolios. Nonetheless, these findings confirm our conclusion that the TDA-based selection strategy, in particular, based on kernel $K_{2}$, outperforms the benchmark models, thus providing an excellent asset selection criteria. [^19] We also observe that the $MVA$ portfolio has a higher out-of-sample mean, SR, and CEQ compared to those from the naïve $1/N$ portfolio. We also perform hypothesis tests to assess whether the values from $MVA$ and naïve strategy are statistically significant. We observed that the mean and CEQ from $MVA$ is significantly greater than that from naïve strategy at 99% confidence level while the Sharpe ratios from the two are not significantly different, a similar conclusion was drawn in DeMiguel et al. (2009b). Panel B in Table 9 shows the out-of-sample performance of the TDA-based sparse global minimum variance portfolios. A similar conclusion can be drawn that in a $K_{2}$ based portfolio, $GV2$ outperforms other TDA-based portfolios, with the exception that $GV5$ has a lower turnover ratio than $GV2$, but the difference is not very large. Further, $GV2$ beats the benchmark portfolios $GVA$, the naïve, and the index in terms of mean and Sharpe ratio. These findings were further confirmed using the statistical tests for mean, SR, and CEQ; for instance, we observed that $MV2$ has a significantly higher mean, SR, and CEQ than $MVA$ portfolios at a 95% confidence level.
Table 9: In this table, we report the out-of-sample mean return, Sharpe ratio (SR), certainty-equivalent (CEQ) return, and portfolio turnover of the optimal mean-variance ($MV$) and global minimum variance ($GMV$) portfolios. The portfolios $MV1-MV6(GMV1-GMV6)$ correspond to the sparse optimal mean-variance (global minimum variance) portfolios obtained based on Strategy 2 corresponding to six similarity matrices. The portfolios $MVA$ and $GMVA$ correspond to the optimal mean-variance and global minimum variance portfolios obtained on all assets, and the naive corresponds to an equal-weighted portfolio. We calculate these measures using realized simple returns of the corresponding portfolios over the out-of-sample period. We also report these measures for the index for comparison. The best and second-best values are highlighted in bold and italics, respectively.
A. Mean-Variance Portfolios
MV1 MV2 MV3 MV4 MV5 MV6 MVA Naïve Index
Mean 9.69E-04 1.82E-03 1.22E-03 7.67E-04 7.27E-04 8.83E-04 1.70E-03 5.89E-04 4.70E-04
SR 4.67E-02 7.33E-02 5.58E-02 3.28E-02 3.55E-02 4.19E-02 6.88E-02 6.22E-02 5.21E-02
CEQ 7.54E-04 1.51E-03 9.79E-04 4.94E-04 5.17E-04 6.61E-04 1.40E-03 5.44E-04 4.29E-04
TR 1.74E+00 1.37E+00 1.78E+00 1.67E+00 1.45E+00 1.39E+00
B. Global Minimum Variance Portfolios
GV1 GV2 GV3 GV4 GV5 GV6 GVA Naïve Index
Mean 3.50E-04 4.42E-04 3.91E-04 3.44E-04 3.95E-04 3.35E-04 3.96E-04 5.89E-04 4.70E-04
SR 5.04E-02 6.45E-02 5.75E-02 4.93E-02 5.73E-02 4.84E-02 6.38E-02 6.22E-02 5.21E-02
CEQ 3.26E-04 4.18E-04 3.68E-04 3.20E-04 3.72E-04 3.11E-04 3.77E-04 5.44E-04 4.29E-04
TR 1.46E+00 1.24E+00 1.70E+00 1.67E+00 1.22E+00 1.26E+00
To check our strategy’s robustness, we study the in-sample and out-of-sample statistics by calculating the mean, SR, and CEQ values for each window (in-sample and out-of-sample) and then taking the average across the windows. Tables 10 presents the in-sample and out-of-sample statistics and their differences. We observe a performance degradation in the out-of-sample period, which is intuitive since the portfolios are optimized over the in-sample period and are tested in the out-of-sample period where no optimization occurs. Further, for $MV2$ and $GV2$, the difference between the in-sample and out-of-sample average values is less than the $MVA$ and $GVA$ portfolios, which suggests that TDA-based filtering strategy is more robust to the changing market dynamics.
Table 10: Average (calculated across the windows) in-sample and out-of-sample values of Mean, SR and CEQ for the portfolios MV1-MV6 and GV1-GV6. The difference in the in-sample and out-of-sample values acts as an indicator for consistency and robustness of the portfolios. Minimum value of the difference is highlighted in bold
A. Mean-Variance Portfolios
MV1 MV2 MV3 MV4 MV5 MV6 MVA
Mean In 3.521E-03 4.612E-03 3.825E-03 3.878E-03 3.742E-03 3.750E-03 5.530E-03
Out 9.689E-04 1.825E-03 1.217E-03 7.668E-04 7.269E-04 8.835E-04 1.704E-03
Difference 2.552E-03 2.787E-03 2.608E-03 3.111E-03 3.015E-03 2.867E-03 3.826E-03
SR In 1.846E-01 1.806E-01 1.921E-01 1.711E-01 1.853E-01 1.880E-01 2.186E-01
Out 7.870E-02 8.542E-02 7.898E-02 4.842E-02 5.695E-02 7.132E-02 9.374E-02
Difference 1.059E-01 9.519E-02 1.131E-01 1.227E-01 1.283E-01 1.167E-01 1.248E-01
CEQ In 3.310E-03 4.242E-03 3.592E-03 3.561E-03 3.498E-03 3.516E-03 5.158E-03
Out 7.517E-04 1.513E-03 9.741E-04 4.901E-04 5.137E-04 6.597E-04 1.394E-03
Difference 2.558E-03 2.729E-03 2.618E-03 3.071E-03 2.984E-03 2.856E-03 3.765E-03
B. Global Minimum Variance Portfolios
GV1 GV2 GV3 GV4 GV5 GV6 GVA
Mean In 5.137E-04 5.227E-04 5.067E-04 5.289E-04 5.276E-04 4.861E-04 5.017E-04
Out 3.504E-04 4.419E-04 3.911E-04 3.443E-04 3.955E-04 3.355E-04 3.961E-04
Difference 1.633E-04 8.084E-05 1.156E-04 1.846E-04 1.321E-04 1.506E-04 1.055E-04
SR In 1.009E-01 1.044E-01 1.040E-01 9.789E-02 1.041E-01 9.965E-02 1.274E-01
Out 7.462E-02 9.952E-02 8.179E-02 7.416E-02 8.972E-02 8.321E-02 9.334E-02
Difference 2.628E-02 4.830E-03 2.219E-02 2.373E-02 1.437E-02 1.643E-02 3.409E-02
CEQ In 4.961E-04 5.052E-04 4.902E-04 5.092E-04 5.102E-04 4.689E-04 4.907E-04
Out 3.260E-04 4.181E-04 3.676E-04 3.195E-04 3.714E-04 3.111E-04 3.765E-04
Difference 1.702E-04 8.713E-05 1.226E-04 1.897E-04 1.388E-04 1.578E-04 1.142E-04
COVID-19 Robustness Check Next, we analyze the portfolio performance in the more volatile period from July 2019 to mid-November 2020, which includes the COVID-19 pandemic. Table 11 presents the out-of-sample performance. Panel A clearly shows that not all sparse TDA-based portfolios perform well. This observation is not entirely surprising because of the increased market fluctuations during the pandemic. Nevertheless, the TDA-based portfolio $MV_{2}$ outperforms the $MVA$, the naive portfolio, and the benchmark index. This is demonstrated with significant gaps; for instance, the Sharpe ratio of 4.63E-02 achieved by $MV_{2}$ is significantly better (almost double) than 2.24E-02 and 2.86E-02 achieved by $MVA$ and the naive portfolio, respectively. Unlike TDA-based mean-variance portfolios, there is no clear pattern in the performance of TDA-based global minimum variance portfolios. Nonetheless, these portfolios, in particular, $GV_{2}$ outperform $GVA$ and naïve portfolio, thus confirming our findings that TDA-based portfolios, in particular, based on the kernel $K_{2}$ are not only sparse but also perform superior to traditional models even in turbulent market conditions, safeguarding against significant losses. These results underline the out-of-sample efficacy of TDA-based models in managing risk and delivering returns.
Table 11: Out-of-sample performance analysis of portfolios $MV1-MV6,MVA,GV1-GV6,GVA$, naïve strategy and the index on COVID-19 data set. The best and second-best values are highlighted in bold and italics, respectively.
A. Mean-Variance Portfolios
MV1 MV2 MV3 MV4 MV5 MV6 MVA Naïve Index
Mean -4.22E-04 1.53E-03 -1.64E-03 1.08E-04 6.09E-04 -8.28E-04 7.66E-04 6.14E-04 6.73E-04
SR -1.60E-02 4.63E-02 -4.81E-02 3.57E-03 1.85E-02 -2.64E-02 2.24E-02 2.86E-02 3.49E-02
CEQ -7.67E-04 9.82E-04 -2.22E-03 -3.52E-04 6.81E-05 -1.32E-03 1.81E-04 3.84E-04 4.87E-04
TR 2.00E+00 1.81E+00 1.86E+00 1.82E+00 1.91E+00 1.56E+00
B. Global Minimum Variance Portfolios
GV1 GV2 GV3 GV4 GV5 GV6 GVA Naïve Index
Mean 7.40E-04 5.68E-04 6.27E-04 8.25E-04 3.86E-04 2.90E-05 1.78E-04 6.14E-04 6.73E-04
SR 6.16E-02 3.01E-02 5.25E-02 4.73E-02 2.12E-02 1.65E-03 1.53E-02 2.86E-02 3.49E-02
CEQ 6.68E-04 3.90E-04 5.55E-04 6.73E-04 2.20E-04 -1.25E-04 1.10E-04 3.84E-04 4.87E-04
TR 1.54E+00 1.91E+00 1.47E+00 1.94E+00 1.74E+00 1.64E+00
6 Conclusion
In this paper, we introduced new distance measures on the space of persistence diagrams and landscapes. Notably, these distance measures incorporate the essential time component of a time series- a facet often overlooked by traditional distance measures. Employing these novel measures as an input to the affinity propagation clustering and proposed two strategies for asset selection to promote sparsity in the portfolio. Specifically, we presented a novel, entirely data-driven approach based on topological data analysis for selecting sparse portfolios in the context of index tracking and Markowitz portfolios. Our empirical findings demonstrate that the performance of sparse portfolios selected based on proposed novel distance measures is not merely comparable, but often excels in volatile market conditions relative to the prevailing methods across diverse set of performance measures. Thus, our paper contributes to sparse portfolio selection and applications of topological data analysis in finance.
Appendix
6.1 Case study: Time series clustering
In this section, we demonstrate the superior performance of the newly introduced distance measures by applying clustering on a standard data set from the machine learning literature. More precisely, we show that the distance measures $AWD$ and $DWD$ that quantify the dissimilarity between persistence diagrams ${\mathcal{D}}_{x}$ and ${\mathcal{D}}_{y}$ outperform the conventional Wasserstein distance ($WD$). Furthermore, we show that the proposed distance measure on landscapes $ALD$ and $DLD$ outperform the traditional distance between landscapes ($LD$), defined as $||\lambda_{x}-\lambda_{y}||_{p}$ where $\lambda_{x}$ and $\lambda_{y}$ are landscapes corresponding to the PD ${\mathcal{D}}_{x}$ and ${\mathcal{D}}_{y}$. For comparison purposes, we also compare the traditional distances, denoted by $d_{1}$ and $d_{2}$, defined in terms of Spearman’s rank correlation $\rho_{1}$ and Pearson’s correlation coefficient $\rho_{2}$, as follows $d_{i}(x,y)=\sqrt{2(1-\rho_{i}(x,y))}~{}~{}i\in\{1,2\}.$ Finally, for completeness, we also consider the Euclidean distance $ED$. We evaluate the clustering performance in terms of the accuracy ($ACC$) He et al. (2019), which ranges from 0 to 1, with a higher value indicating better clustering results.
We collect the data set from the UCI machine learning repository that contains 600 examples of control charts with 60 instances each, synthetically generated by the process in Alcock et al. (1999). The data consists of six different classes of control charts: Normal and cyclic, increasing and decreasing trend, upward and downward shift.
Table 12: Accuracy of clustering results with different distance metrics. Each value $v$ is to be read as $v\times 10^{-2}$
ACC
Distance metric K-med APC with Gaussian kernel
WD 48.3 55
LD 36.67 28.3
AWD 60 56.67
ALD 35 45
DWD 60 61.67
DLD 31.67 33.33
$d_{1}$ 33.3 15
$d_{2}$ 31.67 15
ED 51.67 15
Table 12 presents the clustering results based on two standard algorithms, the affinity propagation clustering (APC) with $K$ clusters and the $K$-medoids with the distance measures mentioned above. [^20] We use Gaussian exponential kernel as similarity measure defined using these distance measures as an input for APC. We fix embedding dimension $d=2$, and we take $\tau=1$ with $p=1$ for calculating the TDA-based distances. Moreover, we focus on the 1-dimensional features, i.e., loops. Further, the hyper-parameter $\sigma^{2}$ in the kernels is taken to be .01. We set $K$ equal to six for illustration purposes. We observe that the TDA-based distance measures AWD and DWD outperform the other distance measures, particularly the existing Wasserstein distance. We also compare the accuracy values reported for the same data set in two recent papers Cheng et al. (2020) and He et al. (2019). The authors proposed a new clustering algorithm and a more accurate distance metric in both papers and compared the performance with different clustering methods. The accuracy values obtained using TDA-based distance measures DWD with affinity propagation clustering and AWD and DWD with K-medoids clustering are higher than the accuracy values reported in Cheng et al. (2020) which is .58 and in He et al. (2019) which is .5886.
6.2 Analyzing the assets selected for index tracking
It is essential to understand the characteristics of assets that are selected based on the similarity matrices $K_{1}-K_{7}$ for the index tracking problem. Specifically, we focus our study on the standard deviation, VaR, and CVaR of the difference return series, that is, the asset return series minus the index return series. Additionally, we analyze downside deviation from the index, and the correlation with the index. [^21] VaR of the differenced series is the VaR relative to the index, which is the largest value by which the portfolio return can miss the index target in $1-\alpha$ fraction of cases. Similarly, CVaR relative to an index shows the mean deviation relative to the benchmark in the worst $\alpha$ cases. Our analysis focuses on the top 60 stocks, identified as the most similar to the index, in each of the 119 windows for all seven similarity matrices. The heat maps in Figures (6(a)-6(e)) illustrate the average values of these measures across all windows, providing insights into the properties of the shortlisted stocks.
There are several observations. First, the assets selected using the similarity matrices $K_{6}$ and $K_{5}$ have a higher correlation than TDA-based similarity matrices, with $K_{6}$ achieving the highest correlation, which is not surprising given that the distance measure in this case is based on Pearson’s correlation. On the other hand, among TDA-based similarity matrices, $K_{2}$ and $K_{4}$ filter assets have a higher correlation with the index than $K_{1}$, $K_{3}$. Second, in terms of standard deviation and downside deviation from the index, the similarity matrices $K_{2}$ and $K_{7}$ have the lowest values among all kernels. Third, the assets filtered using the similarity matrices $K_{2}$ and $K_{7}$ have the lowest values of the VaR and CVaR of the difference series (assets returns and index returns). Other kernels show varying behavior in different windows, but on average, assets picked using $K_{3}$, $K_{5}$, and $K_{6}$ have a higher tail deviation from the index. In conclusion, the assets selected using $K_{2}$ are most suitable for the index tracking problem, as further confirmed in Section 5. The robustness of the kernel $K_{2}$ is established from Figures 7(a)-7(e), which shows similar plots but for a randomly chosen window during the COVID-19 period.
(a) Correlation with index
(b) Standard deviation of difference series
(c) Downside deviation of difference series
(d) VaR of difference series
(e) CVaR of difference series
Figure 6: Average values of five measures (correlation, standard deviation, downside deviation, VaR, CVaR) across all windows for the top 60 stocks (rows) and for the similarity matrices $K_{1}-K_{7}$ (columns)
(a) Correlation with index
(b) Standard deviation of difference series
(c) Downside deviation of difference series
(d) VaR of difference series
(e) CVaR of difference series
Figure 7: Values of five measures (correlation, standard deviation, downside deviation, VaR, CVaR) for a randomly selected window during the COVID-19 period. The values are plotted for the top 60 stocks (rows) and for the similarity matrices $K_{1}-K_{7}$ (columns)